### v6

# Who is Responsible? Data, Models, or Regulations, A Comprehensive Survey on Responsible Generative AI for a Sustainable Future

Abstract.
Generative AI is rapidly moving from research to deployment, elevating the need for responsible development, evaluation, and governance. We conduct a PRISMA-guided review of 232 studies (November 2022–December 2025), spanning large language models, vision-language models, diffusion models, and agentic pipelines. We make four contributions: (1) the first survey bridging governance principles, technical evaluation, and domain deployment across all four system types; (2) a ten-criterion rubric (C1–C10) scoring major AI safety benchmarks on risk-surface coverage, paired with a policy crosswalk mapping benchmarks to regulatory requirements; (3) twelve lifecycle KPIs, explainability guidance for foundation models, and a testbed catalogue; and (4) domain-specific analysis across healthcare, finance, education, arts, agriculture, and defense. Three findings emerge: benchmark coverage is dense for bias and toxicity but sparse for privacy, provenance, deepfakes, and system-level failures in agentic settings; evaluations remain largely static and task-local, limiting audit portability; and inconsistent documentation complicates cross-release comparison. We outline a research agenda prioritizing adaptive multimodal evaluation, privacy and provenance testing, deepfake risk assessment, calibration reporting, versioned artifacts, and continuous monitoring. This survey offers a structured path to align generative AI evaluation with governance needs for safe and accountable deployment.
Project page: Project page
responsible AI, generative AI, AI ethics, bias, privacy, trust, safety
$c$s: Social and professional topics Computing / technology policyccs: Computing methodologies Artificial intelligenceccs: General and reference Surveys and overviews
1. Introduction
Generative AI (GenAI) systems, including large language models (LLMs), vision language models (VLMs), diffusion models, and emerging agentic pipelines, are rapidly transitioning from research prototypes to production deployments that generate text, images, code, and autonomous actions across healthcare, finance, education, and other sectors (Hadi et al., 2023; Hajkowicz et al., 2023). This accelerated deployment heightens the need for responsible development practices that address GenAI-specific risks, including hallucination and unverifiability, prompt injection and jailbreak vulnerabilities, data memorization and leakage, deepfake and media integrity threats, and system-level failures in tool-augmented settings.
Responsible AI (RAI) (Stryker, 2024) has emerged as a central area of research and practice to address these challenges. RAI frameworks emphasize fairness, bias mitigation, privacy protection, security, and the safeguarding of human rights (Peters et al., 2020). By embedding these principles throughout the AI lifecycle, including design, training, deployment, and governance, RAI seeks to realize the societal benefits of AI while minimizing potential harms. While these principles were initially developed for predictive systems, generative models introduce qualitatively new challenges: outputs are open-ended and difficult to enumerate, stochastic decoding produces non-deterministic behaviour, multimodal synthesis creates cross-modal failure modes, and agentic tool use extends the attack surface beyond the model itself. Despite rapid adoption, with 74% of organizations reporting the use of GenAI, only 26% report having a comprehensive organization-wide responsible AI strategy (Capgemini Research Institute, 2024). This imbalance underscores the need to align the development and deployment of these systems with ethical standards, fairness, privacy, security, and societal values.
There has been substantial progress on RAI from both industry (Google AI, 2024; Stryker, 2024; Microsoft, 2024; Anthropic, 2024; Jiang et al., 2025) and academia (Lu et al., 2024b; Dignum and Confidential, 2019; Schiff et al., 2020; Huang et al., 2025). Nevertheless, notable gaps remain in alignment with human values and in operational practice. The academic literature often separates conceptual frameworks from practical implementations, as summarized in Table 1. Conceptual work provides principles and guidelines, whereas applied studies demonstrate deployment in specific contexts. The recent MIT AI Risk Repository analysis (Slattery et al., 2024) indicates that current AI safety frameworks, including those referenced in Table 5, cover only about 34% of risks identified by governance bodies such as the NIST AI Risk Management Framework (of Standards and, NIST), the EU AI Act (Outeda, 2024), and ISO/IEC 42001 (ISO/IEC, 2021b). This gap between governance coverage and practical implementation highlights the need for an integrated review that connects frameworks to implementation strategies and to evaluation methods, including explainable AI and AI-ready testbeds.
In this study, we conduct a comprehensive literature review of responsible GenAI (ResGenAI) frameworks, governance guidelines, evaluation benchmarks, and operational risk assessment methodologies. Our scope centres on the application of RAI principles to generative systems, including LLMs, VLMs, diffusion models, and agentic/RAG pipelines. While we draw on the broader RAI literature for foundational constructs (e.g., fairness, privacy, safety, transparency, accountability), our analysis targets ResGenAI specific risk mechanisms such as hallucination and unverifiability, prompt injections, jailbreaking and tool-use failures, data memorization and leakage, provenance and deepfakes or media integrity risks, intellectual property concerns, and system-level failures in GenAI applications. Despite substantial prior work, ResGenAI guidance remains fragmented across principles, processes, and tools (Table 1). Consequently, there is a pressing need to translate regulatory obligations into concrete engineering controls and to produce portable audit evidence for post-market monitoring.
{forest}
Figure 1. Responsible GenAI landscape organised into three pillars: Challenges (Section 3), Strategies (Section 4), and Applications (Section 5). Technical challenges span data, model/system, and emergent-threat layers; strategies encompass benchmarks, KPIs, explainability, testbeds, governance, and toolkits; applications map domain-specific risks and controls across six sectors.
Contributions. Our key contributions in this work are:
  1. (1)
    To the best of our knowledge, this is the first review article that simultaneously spans governance principles, technical evaluation, and domain deployment for GenAI. We synthesize findings across the NIST AI RMF, EU AI Act, ISO/IEC 42001, and other instruments listed in Table 6, covering LLMs, VLMs, diffusion models, and RAG topics. As shown in Table 1, no prior survey bridges all four thematic areas (CS theory, application, policy, GenAI) and all four novelty dimensions (taxonomy, scope, evaluation, methods) simultaneously.
  2. (2)
    We introduce a ten-criterion rubric (C1–C10) that scores widely used AI safety benchmarks on risk-surface coverage and assurance quality, and a policy crosswalk (Section 4.2.1) that maps each benchmark to the governance requirements it can plausibly evidence. The closest efforts (Liu et al., 2025b; Ferdaus et al., 2026) partially address evaluation taxonomy or trust assessment, but neither links benchmarks to regulatory audit-evidence needs.
  3. (3)
    We operationalise the rubric findings into measurable KPIs (Section 4.2.4), explainability guidance tailored to foundation models (Section 4.3), and a testbed catalogue for continuous evaluation. Section 4.2 links these KPIs to specific benchmarks and audit toolkits, providing practitioners with a decision-support matrix for assembling governance-ready evidence.
  4. (4)
    We apply the RAI lens to real-world applications and domains (Section 5, demonstrating how technical advancements can be effectively aligned with governance frameworks and regulatory requirements.
Table 1. Comparison of related surveys by venue, year, topic, thematic coverage, and novelty dimensions. Checkmarks (✓) indicate coverage; (x) indicates no explicit coverage; (★) indicates partial or limited treatment. Novelty dimensions: Taxonomy = structured risk taxonomy or classification; Scope = multimodal + agentic coverage; Evaluation = governance-aligned benchmarking/rubric; Methods = actionable engineering controls or lifecycle KPIs
Paper Venue Year Topic Thematic Coverage Novelty Dimensions
CS Theory
Application
Policy
GenAI
Taxonomy
Scope
Evaluation
Methods
(Schiff et al., 2020) arXiv 2020 Responsible AI x x x x x
(Mehrabi et al., 2021a) ACM CSUR 2021 Bias and Fairness x x x x x x
(Hu et al., 2021) ACM CSUR 2021 AI Safety x x x x x x
(Trocin et al., 2023) Springer ISF 2021 Responsible AI x x x x x
(Minh et al., 2022) Springer AI Rev. 2022 Explainable AI x x x x x
(Huang et al., 2024a) Springer AI Rev. 2023 Safety x x x x
(Saeed and Omlin, 2023) Elsevier KBS 2023 Explainable AI x x x x x x
(Dwivedi et al., 2023) ACM CSUR 2023 Explainable AI x x x x x x
(Nauta et al., 2023) ACM CSUR 2023 Bias and Fairness x x x x x
(Huang et al., 2023a) IEEE TAI 2023 AI Ethics x x x x x
(Díaz-Rodríguez et al., 2023) Elsevier Inf. Fusion 2023 Responsible AI x x x
(Mhlanga, 2023) Springer Nature 2023 Responsible AI x x x x
(Lu et al., 2024b) ACM CSUR 2024 Responsible AI x x x x x
(Leschanowsky et al., 2024) Elsevier CHB 2024 Privacy and Safety x x x x x
(Caton and Haas, 2024) ACM CSUR 2024 Bias and Fairness x x x x x
(Li and Goel, 2024) Springer ISF 2024 Accountability x x x x x
(Slattery et al., 2024) arXiv 2024 Responsible AI x x x x x x
(Jedličková, 2024) Springer AI & Soc. 2024 Responsible AI x x x x x x
(Al-kfairy et al., 2024) MDPI Informatics 2024 AI Ethics x x x x x
(Sadek et al., 2025) AI and Society 2024 RAI in Practice x x x x x
(Woodgate and Ajmeri, 2024) ACM CSUR 2024 Responsible AI x x x x x
(Reuel et al., 2025) ACM FAccT 2025 RAI (Global maturity) x x
(Gadekallu et al., 2025) arXiv 2025 Best Practices of RAI x x x x x
(Byrne, 2025) ASEE Peer 2025 U.S. RAI Policy x x x x x x x
(Sadek et al., 2025) Springer AI & Soc. 2025 Responsible AI x x x x x
(Baldassarre et al., 2025) IEEE TAI 2025 Responsible AI x x x x x
(Ivanov, 2025) Serv. Ind. J. 2025 RAI in Social Sci. x x x x x x
(Yu et al., 2025) ACM KDD 2025 Trustworthy LLM Agents x x x
(Dong et al., 2025) Springer AI Rev. 2025 LLM Safeguards x x
(Ferdaus et al., 2026) ACM CSUR 2025 Trustworthy LLMs x x x x
(Wang et al., 2025) arXiv 2025 Reasoning Model Safety x x x
(He et al., 2025b) ACM CSUR 2025 LLM Agent Security x x x x
(Papagiannidis et al., 2025) Elsevier JSIS 2025 RAI Governance x x x x x
(Liu et al., 2025b) arXiv 2025 LLM Safety Evaluation x x x x
Ours ACM CSUR 2026 Responsible GenAI
Column definitions CS Theory: conceptual frameworks, taxonomies, or algorithmic contributions. Application: domain-specific deployment or case studies. Policy: governance instruments, regulatory analysis, or compliance guidance. GenAI: explicit focus on generative models (LLMs, VLMs, diffusion, agentic). Taxonomy: structured risk classification beyond a flat listing of concerns. Scope: coverage of multimodal systems and agentic/tool-using pipelines. Evaluation: governance-aligned benchmarking with a scoring rubric or crosswalk. Methods: actionable engineering controls, lifecycle KPIs, or testbed guidance. ★ = partial or limited treatment.
The rest of this paper is organized as follows. Section 2 presents the literature review methodology. Section 3 provides background on ResGenAI. Section 4 reports our results, including thematic findings, benchmark assessments, key performance indicators, explainable AI, and AI-ready testbeds. Section 5 discusses ResGenAI applications, challenges, and strategies. Section 6 presents discussion and forward thoughts, and finally, Section 7 concludes with open challenges and future research directions. The taxonomy presented in Figure 1 provides a structural blueprint for this survey, establishing the thematic connections between high-level technical challenges and domain-specific applications.
2. Methodology
This review adheres to the PRISMA guidelines (Page et al., 2021) to ensure a transparent, reproducible, and rigorous synthesis of evidence. This study follows established methodological recommendations for performing systematic reviews within the computer science community (Al-Selwi et al., 2024; Boland et al., 2017a; Keele et al., 2007). To ensure comprehensive and consistent reporting, the structural framework was modelled after previous extensive surveys in the domain. The mapping procedure utilized in this review is organized into four interconnected phases that covering objectives (Section 2.1), search strategy (Section 2.2), eligibility criteria (Section 2.3), and study selection (Section 2.4) which are visualized in the methodological flow (Fig. 2) and detailed throughout Section 2. For a more exhaustive exploration of these mapping and reporting protocols, foundational works in the field (Boland et al., 2017a; Al-Selwi et al., 2024; Keele et al., 2007; Boland et al., 2017b) offer additional guidance.
Figure 2. The systematic literature mapping process utilized in this study
2.1. Research Questions Formulation
This study conducts a literature review of responsible generative AI (ResGenAI) in the post-GPT period and the synthesis of the findings. We address the following key questions:
  • RQ1: What are the dominant themes, risks, and controls in ResGenAI and multimodal systems since late 2022?
    Motivation: The rapid transition of GenAI from research to production has outpaced traditional risk taxonomies, necessitating a contemporary synthesis of how multimodal and agentic systems introduce qualitatively new failure modes beyond predictive ML.
  • RQ2: Which evaluation facets and metrics are most commonly reported, and where are the gaps?
    Motivation: Identifying the “Evaluation Gap” is critical for determining if current safety suites adequately model the adaptive, iterative pressure applied by real-world adversaries versus merely serving as static, keyword-based refusal scripts.
  • RQ3: What evidence types (empirical, analytical, normative) underpin claims about RAI effectiveness?
    Motivation: A structural deficit in evidence portability hinders institutional trust; this question seeks to evaluate if safety claims are grounded in verifiable technical artifacts or subjective, unvalidated appraisals.
  • RQ4: How can these findings be organized into a practical framework for research and governance?
    Motivation: Principles without practice lead to “safetywashing.” We aim to bridge the gap between regulatory mandates like the EU AI Act and concrete engineering controls, providing an implementation path for verifiable stewardship.
2.2. Evidence Acquisition and Search Strategy
A comprehensive search was conducted across five major academic databases and registers: Scopus ($n=1502$), Web of Science ($n=2798$), IEEE Xplore ($n=279$), ACM Digital Library ($n=656$), and arXiv ($n=78$). To capture the surge in GenAI research following the release of ChatGPT, the search window was limited to publications from November 2022 through December 2025. The search strings utilized an intersection of terms related to Generative AI (e.g., “Large Language Model”, “Foundation Model”, “LLM”) and Responsibility facets (e.g., “AI Safety”, “Trustworthy AI”, “Alignment”, “Hallucination mitigation”). Queries combined RAI concepts (e.g., responsibility, safety, transparency, accountability, fairness, bias mitigation, privacy, equity), with syntax adapted to each database (full strings in Appendix A.1). Syntax was adapted for each database to support field-specific constraints (Title/Abstract/Keywords), and proximity operators were employed where available to enhance precision. To mitigate coverage bias, database queries were complemented with forward and backward snowballing from seed works identified as influential during pilot screening. The review protocol, including sources, screening rules, and the synthesis plan, was prespecified in line with established computer science review guidance (Boland et al., 2017a; Keele et al., 2007)
2.3. Eligibility Criteria
To ensure methodological rigor and relevance to the research questions (RQ1–RQ4), we applied the following criteria:
  • Inclusion Criteria: (1) Peer-reviewed journals, conference proceedings, or high-impact preprints materially shaping RAI discourse; (2) English-language works published within the target 2022–2025 window; and (3) Studies proposing technical frameworks, evaluation metrics, or empirical risk assessments for generative or multimodal systems.
  • Exclusion Criteria: (1) Purely opinionated pieces or editorials without substantive technical argumentation; (2) Short papers or reports lacking sufficient methodological depth; (3) Items whose primary contribution lay outside RAI (e.g., domain applications without safety or governance components); and (4) Redundant duplicates or inaccessible records.
Figure 3. PRISMA flowchart of the systematic literature identification, screening, and inclusion process
2.4. Study Selection and Screening
As detailed in the PRISMA flow diagram (Fig. 3), the initial search identified 5,313 records. Selection was conducted by three independent reviewers to minimize bias. Following automated deduplication ($n=2125$) and the removal of records marked as ineligible by automation tools ($n=609$) or other technical reasons ($n=53$), 2,562 unique records remained for screening.
Titles and abstracts were screened against the eligibility criteria, resulting in the exclusion of 2,153 items. Subsequently, 372 reports were sought for full-text retrieval and assessed for eligibility. Reports were excluded if they failed to meet quality and technical thresholds ($n=140$) or could not be retrieved ($n=37$). This rigorous multi-reviewer process yielded a final synthesis set of 232 studies. Data extraction followed a thematic pipeline to categorize findings into risks, controls, and evaluation facets as defined by the research objectives.
In the following section, we outline the background and key challenges of ResGenAI, which set the stage for the synthesis of our review findings.
3. Background on Responsible Generative AI
3.1. Historical Context
RAI has received a lot of intensified scholarly and policy attention with the rapid application of GenAI (Zirpoli, 2023b; Al-kfairy et al., 2024). GenAI refers to models that synthesize new samples from learned data distributions (van der Zant et al., 2013), spanning language, vision, audio, and multimodal content  (Muneer et al., 2025). Early work combined symbolic rules with probabilistic language models; for example, ELIZA illustrated template driven dialogue in the 1960s (Weizenbaum, 1966). This early era established statistical $n$ gram models (Fink, 2014) and neural sequence modeling for text and speech, laying the groundwork for contemporary large scale generative systems (Qureshi et al., 2025). Figure 4 highlights the most salient milestones for RAI (alignment, evaluation, governance). As shown in the Figure 4. The first major inflection in deep generative modeling arrived with variational autoencoders (VAEs) (Pinheiro Cinelli et al., 2021) and generative adversarial networks (GANs) (Goodfellow et al., 2014), which enabled high-fidelity synthesis and, concurrently, made deepfakes feasible, creating the media integrity challenge that current governance frameworks struggle to address Concurrently, neural sequence-to-sequence models advanced conditional text generation, culminating in the introduction of attention mechanisms and the Transformer (Vaswani et al., 2017), which decoupled sequence modeling from recurrent computation.
A second inflection gave rise to foundation models: large transformers pre-trained in web-scale corpora (e.g., GPT-2/3) and bidirectional encoders (e.g., BERT) that catalyzed transfer across tasks. Alignment advances, including instruction tuning (Peng et al., 2023), reinforcement learning from human feedback (RLHF) (Kaelbling et al., 1996), and constitutional AI (Bai et al., 2022), shifted the emphasis from raw capability to controllability and safety. In vision, diffusion models (Ho et al., 2020) eclipsed GANs for photorealistic generation; when combined with latent compression, they enabled scalable text-to-image synthesis (Rombach et al., 2022).
The most recent phase is multimodal and agentic: vision–language pretraining (e.g., CLIP, Flamingo) (Alayrac et al., 2022), lightweight modality bridging (e.g., BLIP-2) (Li et al., 2023a), and instruction-following multimodal systems (e.g., LLaVA) (Liu et al., 2023b) extend prompting beyond text. Tool use, retrieval-augmented generation (Lewis et al., 2020), and planning/execution patterns operationalize models as agents with external memory and actuation, bringing new safety, privacy, and governance requirements.
Figure 4. Timeline of generative AI milestones
A left-to-right timeline from early rule-based systems to GANs, transformers, and contemporary instruction-tuned large models
3.2. Challenges in Responsible Generative AI
We group GenAI challenges into technical and regulatory domains.
3.2.1. Technical Challenges
Technical issues arise across data pipelines, model design and training, evaluation and assurance, and operational deployment.
Data-Related Challenges
AI models rely on diverse and representative training data to achieve equitable performance across demographic groups and application contexts. Underrepresentation or overrepresentation can induce biased predictions and exacerbate inequities. For GenAI models trained on web-scale corpora, these challenges are amplified by the opacity of crawl pipelines, the infeasibility of manual curation at trillions of tokens, and the risk that synthetic data from prior model generations contaminates future training sets (model collapse). Noisy, incomplete, mislabeled, or stale data also degrade reliability; therefore, adversarial poisoning can seed harmful patterns. Robust validation, anomaly detection, and provenance tracking are critical. Temporal validity is especially challenging in dynamic environments, where outdated data may encode obsolete norms or regulations, undermining relevance and fairness. ResGenAI stewardship, therefore, demands adherence to consent, privacy, and intellectual-property regimes, including emerging AI regulations (Table 5), alongside sector-specific requirements. Datasheets for datasets (Gebru et al., 2021) and data statements (Paullada et al., 2021) document distributions, licensing, collection methods, and intended uses, fostering transparency and accountability.
These data challenges do not operate independently; they interact in ways that current literature undertheorizes. Representativeness goals can conflict with privacy regimes: expanding demographic coverage may require collecting sensitive attributes that GDPR restricts. Quality assurance assumes ground-truth labels, yet for generative training corpora at trillion-token scale, manual verification is infeasible, creating reliance on automated filters (e.g., toxicity classifiers) that themselves embed biases. Most critically, the literature treats these as pre-training concerns, but post-deployment data drift, where user interactions reshape fine-tuning distributions, remains largely unaddressed in the works we reviewed.
Model- and System-Related Challenges
Model- and system-level risks span explainability, bias amplification, robustness under adversarial conditions, calibration, and post-deployment drift. LLMs are often treated as “black boxes,” limiting transparency and accountability in high-stakes settings. These models also inherit and can amplify data biases, producing disparate performance across groups. Aggressive parity constraints can reduce accuracy in safety-critical contexts. In generative settings, these biases manifest not only as allocational harms but as representational harms in generated text and images (e.g., stereotyped depictions, demographic skew, and culturally insensitive outputs), which are harder to enumerate and measure than disparate classification rates. Beyond governance claims, GenAI models also face concrete leakage risks including membership inference and verbatim training-data extraction (Liu et al., 2025a). Technical mitigations such as differential privacy and memorization audits often trade off with model utility (Bashir et al., 2024), necessitating that reporting includes empirical leakage checks rather than relying on governance alignment alone.
Current static, one-shot benchmarks rarely capture these dynamics. Without construct validity and detailed task documentation, headline scores risk falling victim to Goodhart effects (Strathern, 1997a). Safety claims should instead be framed as assurance cases with traceable evidence (e.g., coverage matrices and governance crosswalks), which is a gap we address with our rubric in Section 4.2.1.
3.2.2. Taxonomy of GenAI Hallucinations
A pervasive challenge threatening the reliable deployment of Generative AI across high stakes domains is the phenomenon of hallucination. This refers to the generation of content that appears structurally fluent and plausible yet is factually incorrect, nonsensical, or unfaithful to source premises (Mamo et al., 2023). In the context of robust evaluation and governance, treating hallucinations as a monolithic error class is insufficient for effective risk management. They often stem from complex conflicts between a model’s parametric knowledge learned during pretraining and the provided contextual knowledge in prompts or RAG systems.
To establish a systematic framework for evaluating misinformation specific to ResGenAI, we categorize hallucinations into a five part taxonomy based on technical and contextual failure modes (Banerjee et al., 2024). This taxonomy, visualized in Fig. 5 clarifies the unverifiability risk that current benchmarks in Section 4.2.1 and sectoral applications in Section 5 must address. By mapping these types to underlying technical failure mechanisms and providing illustrative actual impacts, this framework enables practitioners to select appropriate detection metrics and mitigation techniques tailored to the specific nature of the generated error.
Figure 5. A comprehensive taxonomy illustrating five distinct types of generative AI hallucinations, detailing their underlying technical failure mechanisms and providing illustrative real world impacts and cases.
3.2.3. Regulatory and Broader Sociotechnical Challenges
Regulatory challenges concern the design, implementation, and enforcement of frameworks that make GenAI ethical, accountable, privacy-preserving, and safe across domains and jurisdictions. User behavior and deployment context mediate how risks materialize.
The EU AI Act proposes risk tiering and obligations for robustness, transparency, logging, and post-market monitoring (Outeda, 2024). Provenance transparency, especially for copyrighted or sensitive data, is critical. Legal and sectoral expectations for interpretability in high-risk settings meet practical limits as models scale, creating tension between compliance and performance (Slattery et al., 2024); dynamic risks such as election misinformation and deepfakes further require evolving standards. Requirements for meaningful human oversight, user recourse, and disability access are unevenly specified. Over-reliance on assistive GenAI can mask failure modes; documenting escalation paths to human review and accessible interaction patterns is essential in high-stakes contexts. Divergent regimes, such as risk-management frameworks, sectoral law, and emerging standards, create fragmentation across jurisdictions. Conformity assessment, post-market monitoring, and enforcement capacity vary widely, and evidence portability remains an open problem for cross-border deployments. Inconsistent processes for dataset governance, evaluation, deployment, and monitoring further exacerbate bias and privacy risks (Slattery et al., 2024); pragmatic, auditable codes of practice can harmonize controls and enable proportionate assurance across sectors.
3.3. Emergent Threats
While Section 3.2 details established technical risks, we identify further ResGenAI emergent threats that represent a second generation risk landscape. These threats arise from the intersection of increasing model scale, autonomous capabilities, and decentralized deployment patterns.
3.3.1. Adversarial Technical Evolution
Recent evidence suggests that text embeddings can be used to reconstruct original, sensitive training data through representation inversion (Liu et al., 2025a). Unlike direct data extraction, these attacks exploit the vector space to recover high fidelity personal data, rendering traditional anonymization insufficient for models with high memorization rates. Concurrently, the landscape of jailbreaking is shifting from manual roleplay to automated, optimization driven techniques. Using gradient based optimization, adversaries can craft “universal” adversarial suffixes that destabilize safety layers at speed, often achieving success rates that manual engineering cannot reach (Samvelyan et al., 2024).
3.3.2. Sociotechnical and Ecosystem Risks
The rise of decentralized, open source deployments introduces shadow AI risks, where models are fine tuned to remove safety guardrails outside of institutional oversight. Furthermore, we identify the emergence of synthetic identity swarms that are coordinated networks of generative agents that create synthetic consensus cascades to shift human beliefs at an unprecedented scale (Lucas et al., 2023). These threats are compounded by broader sociotechnical concerns: training and serving large models incur material energy and carbon costs with equity implications for global research participation (Khan et al., 2025), while regulatory regimes struggle to anticipate dual use risks without stifling innovation (Weidinger et al., 2022). This complexity creates a “hydra effect”, where mitigating one failure mode, such as toxicity, may inadvertently expose others, including increased sycophancy or declining creative reasoning (Bai et al., 2022). While techniques like RLHF (Kaelbling et al., 1996) can suppress harmful surface level patterns, they often fail to remove underlying harmful representations.
3.3.3. Agentic Specific Failure Vectors
As GenAI evolves toward agentic workflows, where models autonomously plan, select tools, and execute actions and the attack surface expands beyond linguistic outputs to operational environments (Lu et al., 2024a). Unlike single turn interactions, agentic pipelines introduce multi step dependencies that can amplify failure modes. Table 2 systematizes these vectors to guide system level evaluation protocols.
Table 2. Agentic Specific Failure Vectors: Risks in Autonomous Tool Use and Planning
Failure Vector Technical Mechanism Governance and Security Impact
Tool Misuse Unintended or unauthorized API calls resulting from prompt injection or planning errors (Lu et al., 2024a). Leads to financial loss, unauthorized data modification, or service disruption in production environments.
Recursive Hallucination Compound errors where a hallucination in one step serves as the factual basis for subsequent planning (Alansari and Luqman, 2025). Destabilizes reasoning chains, making final autonomous actions entirely ungrounded.
Authentication Leakage Inadvertent exposure of system keys, tokens, or PII during tool calling or log generation (Barth and De Jong, 2017). Creates critical security vulnerabilities that allow adversaries to hijack agentic permissions.
Plan Hijacking Adversarial manipulation of intermediate Chain of Thought to redirect the goal of the autonomous system (Yu et al., 2025). Undermines human oversight by masking malicious intent within seemingly logical execution steps.
This systemic fragility motivates the use of proportionate, auditable controls and transparent evidence. In Section 4.2, we operationalize this via our comparative rubric and provide a governance crosswalk mapping risks to evaluative evidence and compliance narratives.
4. Results
In this section, we present the results of our structured literature review on ResGenRAI. We first report the thematic findings that emerged from our synthesis. Then, we provide a comparative assessment of widely used AI safety benchmarks against our governance-aligned rubric, followed by an analysis of key performance indicators (KPIs), explainability methods, AI-ready testbeds, and applications across domains. The acronyms used in this work are reported in Appendix Table 1.
Figure 6. Trends in Responsible AI Publications and Keyword Evolution Over Time. (A) Number of publications in RAI by Year. (B) Frequency of keywords by year
Publications Trends in RAI by Year
(a) Top themes
(b) Theme trends
(c) Theme co-occurrence
(d) Themes $\times$ publication type
(e) Top keywords
(f) Publications per year
Figure 7. Thematic analysis overview on RAI literature
4.1. Thematic Findings
Figure 7 summarizes thematic distribution, temporal trends, co-occurrence patterns, publication-type mix, and keyword salience. Figure 6 provides a compact view of annual volume and keyword evolution. Appendix Table 2 list representative studies included in this review. Our thematic synthesis highlights that ResGenRAI scholarship in the post-GPT period clusters around five dominant themes: (i) safety and benchmarking, with a surge of multimodal evaluation suites and red-teaming methods; (ii) fairness, bias, and governance, where debates center on aligning technical audits with emerging laws such as the EU AI Act and NIST AI RMF and others (Table 5); (iii) explainability and transparency, underscoring the tension between interpretability and performance in large foundation models; (iv) privacy and security, particularly membership inference and data extraction risks; and (v) sectoral and sociotechnical applications, spanning healthcare, finance, education, and agriculture. Our temporal analysis shows rapid growth from late 2022 onward, with multimodal and agentic systems gaining salience by 2024–2026. These themes frequently co-occur, for example where explainability is invoked to mitigate fairness concerns, or where privacy and governance frameworks intersect in deployment contexts.
4.2. Benchmarks: Comparative Assessment
This section formalizes our governance-aligned rubric (C1–C10) and applies it to widely used benchmarks; we then compare coverage and link results to audit evidence in Tables 4 and 6.
4.2.1. Comparative Assessment Protocol
Our benchmark assessment follows a structured qualitative expert appraisal with ordinal scoring, a well-established approach for rubric-based evaluation in computer science survey research (Harzing and Alakangas, 2016). It is neither a quantitative meta-analysis (no effect sizes or pooled statistics are computed) nor an unstructured qualitative review. Instead, we employ codebook-guided independent scoring that yields ordinal coverage indicators (0/1/2 per criterion), which are then normalized into a composite index. The resulting scores characterize the breadth of risk-surface coverage and documentation quality across benchmarks (they are not cardinal performance measures and should not be interpreted as rankings of overall safety). This approach is appropriate because the underlying evidence, e.g., benchmark papers, dataset cards, public documentation, and is heterogeneous in format and scope, precluding formal quantitative synthesis.
Rubric design.
We assess widely used AI safety benchmarks against a ten-item rubric aligned to risk families and governance needs: C1: Bias/Discrimination, C2 :Toxicity, C3 : Security/Robustness (including adversarial), C4: Misinformation/Disinformation, C5 : Deepfakes/Media Integrity, C6 : Privacy (membership/extraction), C7 : System-level failure, C8 : Malicious-actor realism (red-team/adaptive), C9 : Metric validity/documentation, and C10: Governance alignment (e.g., NIST AI RMF, EU AI Act). The ten criteria (also given in Appendix A.2) reflect risk families that recur across governance instruments and AI risk repositories.
Scoring procedure and reliability.
Each benchmark was scored $s_{i}(B)\in\{0,1,2\}$ per criterion (0 = no coverage, $1$ = partial, $2$ = strong) by two independent reviewers, following the codebook in Appendix A.2. The codebook specifies per-criterion construct definitions, qualifying evidence types, scoring anchors, and a disambiguation rule that assigns tasks spanning multiple risk families to the primary stated objective to prevent double-counting. Evidence was drawn according to a priority ordering: (i) official paper and documentation; (ii) dataset or model cards and maintained leaderboards; (iii) widely cited third-party reports. Disagreements were resolved through structured discussion and, where necessary, adjudication by a third reviewer. Equal weights are used by default. The normalised score is
(1) $S(B)=\frac{\sum_{i=1}^{10}w_{i}\,s_{i}(B)}{2\,\sum_{i=1}^{10}w_{i}},\quad w_{i}>0.$
This yields values in $[0,1]$, where 0 indicates no coverage and $1$ indicates full coverage across all criteria. To verify that our conclusions are not an artefact of the equal-weight assumption, we conducted a sensitivity analysis under alternate weight profiles that prioritise high-stakes criteria (e.g., C3 Security, C6 Privacy); the resulting Spearman rank correlation exceeded $\rho>0.85$, indicating that comparative breadth is a stable indicator of evaluative maturity.
Table 3. Mapping benchmarks to risk surface areas. $\checkmark$ = supported; $\triangle$ = partial; – = no explicit coverage (from public papers and documentation). This crosswalk expands on rubric criteria C9 (metric validity/documentation) and C10 (governance alignment) from Table 4.
Benchmark Bias/Discrimination Toxicity Security Misinfo Deepfakes Privacy System fail. Malicious
HarmBench (Mazeika et al., 2024) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
SALAD-Bench (Li et al., 2024a) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\triangle$
ALERT (Tedeschi et al., 2024) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Risk Taxonomy (Cui et al., 2024) $\checkmark$ $\checkmark$ $\triangle$ $\triangle$ $\triangle$ $\triangle$
TrustGPT (Huang et al., 2023b) $\checkmark$ $\checkmark$
RealToxicityPrompts (Gehman et al., 2020) $\checkmark$
DecodingTrust (Wang et al., 2023a) $\checkmark$ $\checkmark$ $\checkmark$ $\triangle$ $\triangle$ $\triangle$
SafetyBench (Zhang et al., 2024) $\checkmark$ $\checkmark$ $\triangle$
MM-SafetyBench (Liu et al., 2024) $\checkmark$ $\triangle$ $\triangle$ $\triangle$ $\checkmark$
XSTest (Röttger et al., 2024) $\triangle$ $\checkmark$ $\triangle$
Rainbow Teaming (Samvelyan et al., 2024) $\triangle$ $\checkmark$ $\triangle$ $\checkmark$ $\checkmark$
MLCommons AI Safety v0.5 (Vidgen and et al., 2024) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
HELM (Liang et al., 2023) $\checkmark$ $\checkmark$ $\triangle$ $\checkmark$
SHIELD(Shi et al., 2024) $\triangle$ $\triangle$ $\checkmark$
BIG-bench (bench authors, 2023) $\checkmark$ $\checkmark$ $\triangle$ $\triangle$ $\triangle$
PrivLM-Bench (Li et al., 2024b) $\checkmark$
PrivacyLens (Shao and et al., 2024) $\triangle$ $\checkmark$
HumaniBench (Raza et al., 2025) $\checkmark$ $\triangle$ $\triangle$ $\triangle$ $\triangle$
Table 4. Comparative assessment rubric for AI safety benchmarks (0 = no coverage, 1 = partial, 2 = strong). Scores use equal weights and are normalized to $[0,1]$. Per-criterion scoring anchors and evidence rules are provided in Appendix A.2. This table reports scores across all ten criteria (C1–C10), including C9 (metric validity/documentation) and C10 (governance alignment).
Benchmark C1 C2 C3 C4 C5 C6 C7 C8 C9 C10 Score
RealToxicityPrompts (Gehman et al., 2020) 0 2 0 0 0 0 0 0 1 0 0.15
SafetyBench (Zhang et al., 2024) 1 1 1 0 0 0 0 0 1 1 0.25
MM–SafetyBench (Liu et al., 2024) 1 1 1 0 1 0 0 0 1 1 0.30
SALAD Bench (Li et al., 2024a) 1 1 1 0 0 0 0 0 1 1 0.25
HarmBench (Mazeika et al., 2024) 0 1 1 0 0 0 0 1 1 0 0.20
Rainbow Teaming (Samvelyan et al., 2024) 1 1 2 1 0 0 1 2 1 1 0.50
ALERT (Tedeschi et al., 2024) 0 1 1 0 0 0 0 2 1 1 0.30
HELM (Liang et al., 2023) 1 1 1 1 0 0 0 0 1 1 0.30
PrivLM Bench (Li et al., 2024b) 0 0 0 0 0 2 0 0 1 1 0.20
PrivacyLens (Shao and et al., 2024) 0 0 0 0 0 2 0 0 1 0 0.15
MLCommons AI Safety v0.5 (Vidgen and et al., 2024) 1 1 1 1 0 0 0 1 1 1 0.35
SHIELD (Shi et al., 2024) 0 0 1 0 2 0 0 0 1 0 0.20
MFG for GenAI (Foundation, 2024) 1 0 1 1 0 0 0 1 0 0 0.20
XSTest (Röttger et al., 2024) 0 0 0 0 0 2 0 1 1 0 0.20
DecodingTrust (Wang et al., 2023a) 1 1 1 0 0 1 0 1 1 1 0.35
TrustLLM (Huang et al., 2024b) 1 1 0 1 0 1 1 0 1 0 0.30
BIG bench (bench authors, 2023) 1 1 1 0 0 0 1 0 0 0 0.25
Risk Taxonomy (Cui et al., 2024) 1 1 1 0 0 1 0 1 1 2 0.40
TrustGPT (Huang et al., 2023b) 1 1 0 0 0 0 0 0 1 0 0.20
HumaniBench (Raza et al., 2025) 2 1 0 1 0 1 1 0 1 1 0.40
Criteria (C1 to C10) C1 Bias/Discrimination; C2 Toxicity; C3 Security/Robustness; C4 Misinformation/Disinformation; C5 Deepfakes/Media Integrity; C6 Privacy (membership/extraction); C7 System level failure (tool/agentic); C8 Malicious actor realism (red team/adaptive); C9 Metric validity/documentation; C10 Governance alignment. Normalization (equal weights): $S(B)=\frac{\sum_{i=1}^{10}s_{i}(B)}{20}$ with $s_{i}\in\{0,1,2\}$.
Table 5. Governance-based Frameworks and Technical AI Tools used in RAI Practices
Category Governance-based Frameworks Technical AI Solutions
Fairness & Bias Mitigation ISO/IEC TR 24027:2021 (ISO/IEC, 2021a): technical report on AI and data quality. NIST SP-1270 (Schwartz et al., 2022): framework for identifying and managing bias in AI. AI4ALL (AI4ALL, [n. d.]): framework for increasing diversity & inclusion in AI education, research, & policy. African Union (AU) Data Policy Framework (African Union, 2022): promotes equitable, inclusive AI and data governance across member states. Singapore IMDA & PDPC’s AI Governance Framework (Infocomm Media Development Authority and Personal Data Protection Commission (PDPC)(2020), IMDA): addresses fairness, human-centricity, and non-discrimination in AI decision-making. Google What-If Tool (Wexler et al., 2020): evaluate for model fairness via visualizations. IBM AI Fairness 360 (AIF360) (Bellamy et al., 2019): tools for evaluation and mitigation of bias. Microsoft Fairlearn (Weerts et al., 2023): assess and enhance fairness of ML models.
Interpretability & Explainability IEEE P2976 (Polemi et al., 2024): standards for Explainable Artificial Intelligence. IEEE P2894 (Committee, 2024): guidelines for implementing explainable AI technologies. UK ICO & The Alan Turing Institute: Explaining Decisions Made with AI (Information Commissioner’s Office and The Alan Turing Institute, 2022). China CAC Algorithm Recommendation Provisions (Cyberspace Administration of China, 2022b): mandates transparency and explainability requirements for algorithmic recommendation systems. Learning Interpretability Tool (LIT) (Google PAIR, [n. d.]): model interpretability and explainability. InterpretML (Microsoft Research): toolkit for measuring explainability using LIME and SHAP. IBM AI Explainability 360 (AIX360) (Arya et al., 2019): tools for explainability of ML models.
Transparency & Accountability U.S. Government Accountability Office: AI Accountability Framework (U.S. Government Accountability Office, 2021). OECD AI Principle of Accountability (Organisation for Economic Co-operation and Development, 2019). China Measures for the Management of Generative AI Services (CAC, 2023) (Cyberspace Administration of China, 2023): requires service providers to disclose AI-generated content, maintain logs, and be accountable for model outputs. Brazil’s Proposed AI Act (PL 2338/2023) (Senado Federal do Brasil, 2023): mandates impact assessments, transparency obligations, and accountability measures for high-risk AI systems. SDAIA National AI Ethics Principles (Saudi Data and Artificial Intelligence Authority, 2023a): establishes accountability, auditability, and transparency requirements for AI deployments in Saudi Arabia. Evidently (Evidently AI, [n. d.]): evaluate, test and monitor ML models. MLBench (MLBench Team, [n. d.]): improve model transparency, reproducibility, and robustness. AI Verify (Infocomm Media Development Authority ([n. d.]), IMDA): Singapore’s IMDA-developed AI governance testing framework.
Privacy, Safety & Security NASA (Smith and Kearns, 2020): hazard contribution modes of ML in space applications. IEEE P7009 (Farrell et al., 2021): standards for the fail-safe design of autonomous systems. ISO/IEC TS 27022:2021 (ISO/IEC, 2021b): information security management systems guidelines. NIST AI 100-2 E2023 (Oprea and Vassilev, 2023b): specific cybersecurity practices for AI. UK Government AI Cybersecurity Code of Practice (Department of Science, 2024). MIT AI Risk Repository (Slattery et al., 2024): database documenting AI risks and safety gaps. Google’s Secure AI Framework (SAIF) (Google, 2023). Brazil LGPD (da Republica, Brasil): data protection law governing privacy, data processing, and individual rights in AI-driven systems. Nigeria Data Protection Act (NDPA, 2023) (Agency, 2023): establishes data privacy safeguards for AI systems processing personal data of Nigerian citizens. China Personal Information Protection Law (of China, 2021): comprehensive data privacy regulation for AI systems. UAE Federal Decree-Law No. 45 on Personal Data Protection (Government, 2021): governs data privacy and individual rights in AI-driven data processing within the UAE. Unitary (Ltd., [n. d.]): context-aware toxicity tracker. Adversarial Robustness Toolbox (ART) (Research, [n. d.]). Privacy Meter (Privacy and Lab, [n. d.]): audit privacy of ML models. Google Differential Privacy Libraries (Google, [n. d.]). SecretFlow (Group, [n. d.]): privacy-preserving data analysis and ML toolkit. TensorFlow Privacy (Google, [n. d.]). Betterdata.AI (inc., 2024): PET on synthetic data generation.
Ethical Guidelines & Compliance European General Data Protection Regulation (GDPR) (European Parliament and Council of the European Union, 2016). Canadian AI and Data Act (Innovation, Science and Economic Development Canada, 2023a). Canadian Voluntary Code of Conduct on the Responsible Development and Management of Advanced GenAI Systems (Innovation, Science and Economic Development Canada, 2023b). American Executive Order on Safe, Secure, and Trustworthy AI (House, 2023). China’s New Generation AI Development Plan (AIDP, 2017) & AI Governance Principles (2019) (State Council of the People’s Republic of China, 2017; National New Generation AI Governance Expert Committee (2019), China): outline China’s national strategy for AI development with ethical principles for security, fairness, and human-centric AI. UAE National AI Strategy 2031 (United Arab Emirates Government, [n. d.]): comprehensive national strategy establishing ethical AI goals and responsible AI deployment. SDAIA Responsible AI Guidelines (Saudi Data and Artificial Intelligence Authority, 2023b): ethical guidelines for responsible AI use and regulatory compliance in Saudi Arabia. Singapore Model AI Governance Framework (2nd Ed., IMDA) (Infocomm Media Development Authority and Personal Data Protection Commission (PDPC)(2020), IMDA): practical guidance on responsible AI deployment covering internal governance and human oversight. AU Continental AI Policy Framework (2024) (African Union, 2024): continent-wide framework for ethical, inclusive, and human rights-respecting AI governance. Colombia’s AI Ethics Framework (Ministerio de Tecnologías de la Información y las Comunicaciones (2019), MinTIC) & Chile National AI Policy (Ministerio de Ciencia, Tecnología, Conocimiento e Innovación, 2021): regional strategies for responsible AI adoption. Google Responsible AI Practices (Google, [n. d.]). Australia’s AI Ethics Principles (Australian Government Department of Industry, Science and Resources, [n. d.]). Microsoft’s Responsible AI Toolbox (principles and approach, 2024). OneTrust (OneTrust LLC, [n. d.]): demonstrate trust, measure and manage risk, and compliance. WHO: Ethics and Governance of AI for Health (World Health Organization, 2024). Project Moonshot (AI Verify Foundation, 2024): open-source LLM evaluation and red-teaming toolkit to benchmark GenAI systems against safety and ethical compliance standards.
Monitoring & Auditing AI Incident Database (Responsible AI Collaborative, [n. d.]): index of harms realized in the real world by AI systems. UK ICO AI Auditing Framework (UK Information Commissioner’s Office (2020), ICO). China CAC Measures for the Management of Deep Synthesis Internet Information Services (2022) (Cyberspace Administration of China, 2022a): requires labeling, monitoring, and auditing of AI-generated synthetic media. Smart Africa Alliance AI Audit Guidelines (Smart Africa, 2023): monitoring and audit guidance for AI systems in African public services. IDB AI Governance Toolkit (Inter-American Development Bank (2023), IDB): monitoring and assessment tools for responsible AI deployment across Latin American and Caribbean nations. FairVis (Poloclub, [n. d.]): audit classification models for intersectional bias. Aequitas (Center for Data Science and Public Policy, [n. d.]): bias and fairness audit toolkit. Audit-AI : auditing of bias in generalized ML applications.
Table 6. Governance alignment crosswalk (illustrative). ✓= supported evidence; $\triangle$= partial; –= not covered. A ✓ indicates plausible evidentiary support rather than a statement of legal compliance; see Appendix B for the non-compliance note. This crosswalk expands on rubric criteria C9 (metric validity/documentation) and C10 (governance alignment) from Table 4.
Requirement
RTP
SafetyBench
MM–SafetyBench
SALAD
HarmBench
Rainbow
ALERT
HELM
PrivLM
SHIELD
MLC v0.5
HumaniBench
NIST: Privacy & Security (of Standards and, NIST, NIST) $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$
EU AI Act: Robustness/Resilience  [^1] https://artificialintelligenceact.eu/article/15 $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$
EU AI Act: Transparency/Record-keeping  [^2] https://artificialintelligenceact.eu/article/12/ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$
Multimodal Deepfake Risk (C2PA, 2024) $\triangle$
Adaptive/Red-team Evidence (of Standards and, NIST) $\triangle$ $\triangle$ $\triangle$ $\triangle$
Abbreviations RTP = RealToxicityPrompts; MM–Saf. = MM–SafetyBench; HarmB. = HarmBench; MLC v0.5 = MLCommons AI Safety v0.5; Humani = HumaniBench.
4.2.2. Comparative Assessment Findings.
Building on the challenges in Section 3.2, we apply the rubric in Section 4.2.1 to evaluate benchmarks against governance needs. Table 3 presents the risk-surface matrix for criteria C1–C8. The full rubric scores, including C9 (metric validity and documentation) and C10 (governance alignment), are given in Table 4. Table 6 then provides a finer-grained governance crosswalk mapping each benchmark to specific regulatory requirements.
Risk-surface coverage.
Table 3 presents the representative benchmarks across eight salient risk families (C1–C8). Entries are based on public papers and documentation and reflect explicit task/metric coverage rather than incidental behaviour. The results in Table 3 show that coverage is densest for toxicity and bias (e.g., RealToxicityPrompts, SafetyBench, HarmBench), while privacy is siloed to specialized suites (PrivLM-Bench, PrivacyLens). Deepfake/media integrity is sparsely addressed (notably SHIELD), and only a few efforts incorporate adaptive adversaries or richer system context (Rainbow Teaming, ALERT). System-failure probing is rare, appearing primarily in Rainbow Teaming.
Rubric-scored comparison.
We operationalize the ten-criterion rubric in Section 4.2.1 to produce a normalized $[0,1]$ score per benchmark. The aim is not to rank “best overall,” but to indicate breadth and governance-ready evidence across risk families and evaluation qualities (metric validity, documentation, policy alignment). The results in Table 4 show that Rainbow Teaming attains the highest breadth score (0.50), reflecting strong coverage of security under adaptive attack and red-team realism. Risk Taxonomy and HumaniBench (0.40 each) combine multi-risk awareness with governance-facing structure. MLCommons AI Safety v0.5 and DecodingTrust (0.35) provide broad but mostly static probes. Single-issue suites score lower by design, e.g., RealToxicityPrompts and PrivacyLens (0.15), useful for their focus but limited for governance evidence.
Governance context and crosswalk.
Table 5 situates the analysis within widely used frameworks and tools. Table 6 then maps each benchmark to selected governance requirements and risks such as multimodal deepfake risk, and evidence for adaptive/red-team evaluation. The results show that only a subset of suites provide tangible evidence for privacy/security controls (e.g., PrivLM-Bench). Robustness/resilience support is strongest where adaptive red teaming is integral (Rainbow Teaming, ALERT). Deepfake risk is narrowly covered (SHIELD), suggesting a gap for multimodal governance. Transparency/record-keeping signals are uneven outside of meta-benchmarks (HELM, MLCommons), which can hinder audit trails under the EU AI Act.
Figure 8. Responsible GenAI governance landscape and linkage to our evaluation criteria. Technical factors correspond primarily to rubric items C1–C7 (fairness/bias, toxicity, security/robustness, mis/disinformation, deepfakes/media integrity, privacy, system-level failure). Risk & assurance processes (anticipation, responsiveness, reflexivity, inclusion, knowledge management) support adaptive testing (C8), metric validity/record-keeping (C9) and Legal instruments (e.g., EU AI Act, NIST AI RMF) ground C10 (governance alignment). Sustainable AI highlights environmental/societal/economic impacts that are only weakly covered by current benchmarks, motivating the gaps observed in Tables 46
4.2.3. Proposed Improvements to AI Safety Benchmarks
Based on the benchmark gaps we found in the literature, we suggest constructive opportunities to complement existing benchmarks and to support broader governance use. First,we recommend that benchmark developers encourage diverse and well-documented corpora across regions, languages, and demographics, accompanied by dataset cards or data statements that capture provenance, consent, licensing, and curation criteria to enhance traceability and accountability (Eriksson et al., 2025; Gebru et al., 2021). Equally important is construct clarity: each task and metric should state explicitly what it is intended to measure and, where feasible, be validated with multidisciplinary input to guard against superficial compliance claims (Ren et al., 2024). Without these foundations, high benchmark scores risk measuring test-taking ability rather than genuine safety.
Incorporating adversarial evaluation, including undisclosed items and evolving prompts can be used bring benchmarks in line with established security testing practices (Oprea and Vassilev, 2023b; Ghosh et al., 2025). Scope should also expand to cover multimodal, non-English, and higher-risk domains, as well as multi-agent or autonomous settings where system-level failures are more likely to surface (Salazar et al., 2025). These benchmarks rarely make explicit linkages to policy instruments. Engaging multiple stakeholders and mapping evaluation outputs to governance frameworks would mitigate commercial or domain bias and improve regulatory utility (Outeda, 2024). Continuous evaluation through automated pipelines that enable periodic re-testing across the model lifecycle would further improve comparability over time (Wang et al., 2024). Finally, structured red teaming and stress tests under distribution shift should be integrated to surface emergent behaviours that may elude static suites, drawing on practices such as ALERT and Rainbow Teaming.
A broader concern emerging from this analysis is that many existing benchmarks may facilitate what might be termed “safetywashing”: models can achieve high scores on structured benchmarks such as SafetyBench (Zhang et al., 2024) by learning keyword-driven refusal scripts, while failing under the multi-turn, emotionally complex interactions documented in real-world deployment incidents (Roose, 2023). This pattern underscores the need to move beyond static compliance checks toward evaluation paradigms that capture the adversarial and socially situated nature of human–model interaction.
The governance landscape framework is intended to illustrate that responsibility is not a single algorithmic check, but a multi-dimensional stewardship involving technical, legal, and social factors. Figure 8 provides the foundational logic for our ten-criterion rubric, ensuring that our technical assessment is grounded in real-world governance requirements. To operationalize these improvements and make evidence portable across audits, we now specify lifecycle KPIs that teams can track during development, deployment, and post-market monitoring.
4.2.4. Key Performance Indicators (KPIs) for Responsible GenAI
KPIs provide measurable criteria to assess the performance of AI systems across critical dimensions of responsibility (Mattioli et al., 2024). Unlike ethical principles that are mostly subjective, KPIs translate RAI goals into operational metrics that can be tracked, compared, and audited across the model lifecycle. They bridge governance expectations with technical implementation so that fairness, transparency, robustness, privacy, and sustainability are evidenced through quantifiable measures, enabling compliance and continuous improvement. We adapt established measures (SPD, DI, accuracy parity, robustness under perturbation, and energy) from prior work (Pessach and Shmueli, 2022), and introduce pragmatic, author-defined KPIs for explainability, inclusivity, auditing cadence, and high-stakes errors to support governance-ready evidence.
Notation and conventions $f(x)$ is the model score for input $x$ (state whether probability, logit, or other monotone score); $y\in\{0,1\}$ is the true label; $\hat{y}=\mathbb{1}\{f(x)\geq t\}$ is the predicted label with a single threshold $t$ applied to all subgroups unless stated; $A$ is a protected attribute with subgroups $a$ and $b$; $\phi_{0}$ is a baseline output and $\phi_{i}(x)$ is the attribution for feature $i$ (e.g., SHAP or Integrated Gradients); “accuracy” refers to a stated metric (e.g., top-1, F1, AUROC) computed per subgroup with the same threshold $t$; a perturbation is a noise or attack consistent with a stated threat model; $P_{\mathrm{avg}}$ is average power (W), $T$ is time (h), $\mathrm{CI}$ is grid carbon intensity (kgCO2e/kWh), and $\mathrm{PUE}$ is Power Usage Effectiveness.
  1. (1)
    Data quality and integrity
    $Q=\frac{\mathrm{Valid\ Entries}}{\mathrm{Total\ Entries}}\times 100.$
    This KPI measures the fraction of entries that pass schema, label, and consistency checks (Pessach and Shmueli, 2022). Here, we define “valid” checks to report $Q$ (quality) by subgroup/domain if applicable.
  2. (2)
    Data privacy compliance
    $P=\frac{\mathrm{Compliant\ Data\ Points}}{\mathrm{Total\ Data\ Points}}\times 100.$
    This KPI fractions with documented lawful basis, consent or equivalent, and license compatibility (Leschanowsky et al., 2024). We specify regime (e.g., GDPR/CCPA), provenance fields audited, and audit scope; legal review may be required.
  3. (3)
    Bias detection (statistical parity difference, SPD)
    $\mathrm{SPD}=P(\hat{y}=1\mid A=a)-P(\hat{y}=1\mid A=b).$
[…]
  4. (4)
    Fairness ratio (disparate impact, DI)
    $\mathrm{DI}=\frac{P(\hat{y}=1\mid A=a)}{P(\hat{y}=1\mid A=b)}.$
    This KPI measures the outcome parity as a ratio. It provides guard against division by zero; also report the symmetric form $\min(\mathrm{DI},1/\mathrm{DI})$.
  5. (5)
    Equal performance (accuracy parity)
    $\Delta_{\mathrm{acc}}=\left|\mathrm{Accuracy}_{A=a}-\mathrm{Accuracy}_{A=b}\right|.$
    This measures the absolute gap in a stated accuracy metric across subgroups (Raza et al., 2024). It reports per subgroup and key intersections; keep $t$ fixed across subgroups.
  6. (6)
    Robustness assessment
    $R=\frac{\mathrm{Errors\ under\ perturbation}}{\mathrm{Total\ perturbed\ inputs}}\times 100,\quad\Delta_{\mathrm{acc}}^{\mathrm{rob}}=\mathrm{Acc}_{\mathrm{clean}}-\mathrm{Acc}_{\mathrm{pert}}.$
    It measures the error rate and accuracy drop under stated threat models. We specify attack type and strength (e.g., $\ell_{p}$ bounds), number of steps, and per-attack results. For generative models, robustness is more appropriately measured as attack success rate (ASR) under jailbreak or prompt-injection attempts, safety-refusal consistency under paraphrase, and output stability under temperature and seed variation, rather than misclassification rate under $\ell_{p}$ perturbations
  7. (7)
    Explainability index (author defined)
    $E(x)=\sum_{i=1}^{n}\left|\phi_{i}(x)\right|.$
    This KPI measures the attribution magnitude summary for an instance using SHAP/IG (Lundberg and Lee, 2017; Sundararajan et al., 2017; Vilone and Longo, 2021). We use as a proxy only; consider also reporting attribution concentration (e.g., entropy over $|\phi_{i}|$).
  8. (8)
    Local faithfulness gap (XAI)
    $\mathrm{LFG}(x)=\frac{\left|\,f(x)-\left(\phi_{0}+\sum_{i}\phi_{i}(x)\right)\right|}{\max\left(\left|f(x)\right|,\epsilon\right)}.$
    This KPI measures the relative completeness error of additive explanations (lower is better). We set $\epsilon$ and summarize LFG as a distribution (median, IQR); complements human-readability studies.
  9. (9)
    High stakes error rate (author defined)
    $E_{h}=\frac{\mathrm{Critical\ errors}}{\mathrm{Total\ predictions}}\times 100.$
    This measures incidence of errors that trigger a predefined harm tier (Kunreuther et al., 2002). We define the harm taxonomy and decision costs; consider a harm weighted rate in safety critical settings. In generative deployments, critical errors include hallucinated medical or legal facts, harmful completions, and PII leakage incidents; $E_{h}$ should accordingly be computed as the incidence of such events per $N$ generations in the target domain, with harm severity tiers defined by deployment context
  10. (10)
    Audit frequency and resolution time
    $F_{a}=\frac{\mathrm{Audits}}{\mathrm{Time\ period}},\qquad T_{r}=\mathrm{median\ time\ to\ resolve}.$
    This KPI is cadence of audits and timeliness of remediation (Díaz-Rodríguez et al., 2023; Hasan, 2021). We specify window and scope; report $T_{r}$ with IQR and the share closed within SLA.
  11. (11)
    User inclusivity index (author defined)
    $U=\frac{\mathrm{Inclusive\ features\ implemented}}{\mathrm{Planned\ features}}\times 100.$
    It measures the accessibility and localization coverage (Shams et al., 2023). We enumerate the checklist (e.g., WCAG conformance, screen reader support, localization) and review cadence.
  12. (12)
    Sustainability (energy and emissions)
    $E_{\mathrm{kWh}}=\frac{P_{\mathrm{avg}}}{1000}\times T,\qquad C_{\mathrm{CO_{2}e}}=E_{\mathrm{kWh}}\times\mathrm{PUE}\times\mathrm{CI}.$
    This KPI measures energy use and estimated emissions; pair energy with location specific or market based carbon intensity and data center overhead (Van Wynsberghe, 2021). We state measurement method (e.g., nvidia-smi vs power meter), scope (training or inference), normalization unit (e.g., per 1k tokens), and whether CI is location or market based (e.g., CodeCarbon) (Song et al., 2025).
Sections 56 use this rubric (C1–C10) and the lifecycle KPIs as a running lens to interpret domain evidence and highlight where current practice still lacks governance-ready coverage.
4.3. Explainable AI
Recent GenAI era deployments expose failure modes (hallucination, privacy leakage, deepfakes, and tool-use errors) that cut across domains; we therefore use C1–C10 and the KPI set as a consistent lens throughout the remainder of the survey. Having established how benchmarks cover risk families, we examine explainability methods and what they can and can not do, evidence under audit. Explainable AI (XAI) comprises methods that render model behaviour and outputs intelligible to human stakeholders (Meske et al., 2022). Effective explanations serve multiple roles: they support accountability and trust, enable bias and safety analyses, and facilitate compliance with emerging governance requirements (Baker and Xiang, 2023; Arrieta et al., 2020). In practice, XAI must operate across technical, social, legal, and organisational layers rather than as a purely algorithmic add-on.
Figure 9. Post-hoc explainability within a Responsible AI (RAI) pipeline. A black-box model produces predictions $\hat{y}$ and attribution signals $I(x_{i})$, where $I(x_{i})$ denotes a feature-importance signal (e.g., gradient/saliency or a SHAP value). Interpretability pathways (feature relevance, example-based, surrogates, and visual/text explanations) feed fairness diagnostics (e.g., SPD/DI) and calibration analyses, supporting governance-aligned optimisation $\mathcal{L}_{\mathrm{RAI}}=\mathcal{L}+\lambda\,\mathcal{L}_{\mathrm{fair}}$. For LLMs/VLMs, prompt traces and cross-modal attention provide additional observability
Scope and challenges in the GenAI
Foundation models raise explanation challenges beyond conventional ML. Stochastic generation (e.g., temperature or top-$k$ sampling) produces non-deterministic outputs for identical prompts, complicating reproducibility of explanations (Idé and Abe, 2023). Multimodality couples heterogeneous inputs/outputs (text, image, audio), requiring explanations that account for cross-modal attention and fusion dynamics (Pi, 2023). Hallucinations and brittle prompt sensitivity demand methods that diagnose provenance gaps and reasoning failures. Consequently, GenAI-oriented XAI should expose decoding pathways, visualise modality interactions, and characterise failure modes, not merely attribute tokens.
Methods in practice
XAI methods can be ante-hoc (interpretability built into the model class, e.g., sparse linear models, rule lists, decision trees) or post-hoc (explain an existing black-box locally or globally). Common post-hoc families include feature relevance (saliency/attribution, gradients, SHAP/LIME) for instance-level insight (Lundberg and Lee, 2017; Ribeiro et al., 2016a); example-based explanations (prototypes, nearest neighbours) that ground behaviour in the data manifold (Kenny et al., 2021); surrogate models that approximate decision logic with a transparent model; and visual/textual rationales (heatmaps, attention summaries) for human inspection (He et al., 2023). For LLMs and Vision-language-models (VLMs), these are complemented by prompt-engineering probes, chain-of-thought (CoT) audits, and tool-use traces (Natarajan and Nambiar, 2024). Importantly, attention maps and saliency are not proofs of causal reasoning and should be triangulated with counterfactual tests and ablations (Jain and Wallace, 2019).
Figure 9 situates post-hoc explanations within an RAI pipeline. This framework bridges the gap between abstract interpretability theory and concrete engineering practice. Its specific role is to show how attribution signals and fairness checks (Barocas and Selbst, 2016) can be integrated into a standard pipeline to produce the ”portable evidence” needed for external audits and institutional trust. Predictions $\hat{y}=M_{\phi}(x)$ can be accompanied by gradient-based importance $I(x_{i})=\partial M_{\phi}/\partial x_{i}$ to support local inspection; example-based evidence provides case-level comparators; and visual/text explanations aid end-user understanding. Fairness is evaluated with standard disparity measures such as Statistical Parity Difference (SPD) (Kun, 2015) and Disparate Impact (DI) (Barocas and Selbst, 2016):
$\text{SPD}=P(\hat{y}=1\,|\,A=a)-P(\hat{y}=1\,|\,A=b),\quad\text{DI}=\frac{P(\hat{y}=1\,|\,A=a)}{P(\hat{y}=1\,|\,A=b)}.$
A practical training objective combines task loss with an explicit fairness regulariser,
$\mathcal{L}_{\mathrm{RAI}}=\mathcal{L}(\theta)+\lambda\,\mathcal{L}_{\mathrm{fair}},$
allowing developers to trade off utility and fairness under governance constraints. For trust calibration, we recommend standard metrics such as Expected Calibration Error (ECE) (Pleiss et al., 2017) or the Brier score (Rufibach, 2010); ECE summarises the gap between confidence and empirical accuracy across bins:
$\mathrm{ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{n}\,\big|\mathrm{acc}(B_{m})-\mathrm{conf}(B_{m})\big|.$
Adding canonical sources: ECE and Brier are standard measures (Pleiss et al., 2017). Calibrated probabilities, paired with faithful explanations, yield more reliable user-trust assessments.
A compact theoretical frame
Let $p_{\theta}(\cdot\mid x,c,\omega)$ denote a generative model with input $x$, context $c$ (e.g., system prompt, retrieved documents), and stochastic decoding parameters $\omega$ (temperature, top-$k/p$). The predictive case $f_{\theta}:\mathcal{X}\to\mathcal{Y}$ is recovered as a deterministic special instance
$\mathcal{E}:(x,f_{\theta},\omega)\mapsto z\in\mathcal{Z},$
where $\omega$ indexes stochasticity (decoding temperature, top-$k/p$). Local explanations fix $x$; global explanations summarise behaviour over a distribution on $\mathcal{X}$. For VLMs, $\mathcal{X}=\mathcal{X}_{\text{text}}\times\mathcal{X}_{\text{vision}}$ and $z$ pairs textual rationales with grounded visual regions.
Desiderata (axioms) An explanation $z=\mathcal{E}(x,f_{\theta},\omega)$ is useful when it approximately satisfies: (i) local accuracy (additive decompositions sum to output deviation from a baseline) (Lundberg and Lee, 2017); (ii) sensitivity/monotonicity (if $f_{\theta}$’s dependence on feature $i$ increases at $x$, its attribution should not decrease) (Sundararajan et al., 2017); (iii) invariance to representation-preserving transforms and to behaviourally equivalent models (Sundararajan et al., 2017); (iv) stability (Lipschitz continuity in $x$ and $\omega$(Alvarez-Melis and Jaakkola, 2018); (v) counterfactual consistency (factors highlighted by $z$ admit feasible interventions that flip or materially shift the output); and (vi) parsimony (minimal complexity $\Omega(z)$ subject to fidelity) (Kusner et al., 2017). These properties specialise naturally to multimodal settings via alignment between rationales and grounded visual evidence.
Without committing to a particular algorithm, we report three constructs that make explanations auditable: fidelity via perturbation functionals $\Delta_{f}(x;S)$ on regions or features emphasised by $z$ and compared to size-matched random sets; stability via distances $\delta(\mathcal{E}(x,f_{\theta},\omega),\mathcal{E}(x+\epsilon,f_{\theta},\omega^{\prime}))$ under input noise and seed changes; and governance usefulness by pairing explanations with SPD/DI, calibration (ECE/Brier), and provenance artefacts (prompt/context/tool logs). This elevates explanations from plausible narratives to evidence interfaces suitable for audit.
For foundation models, explanations should expose (i) prompt pathway elements (system/user prompts, retrieved context, tool calls), (ii) decoding state ($\omega$: temperature, top-$k/p$), and (iii) cross-modal grounding (text rationales linked to image regions or clips). These artefacts do not prove causality but enable stability and counterfactual checks and support post-hoc accountability (Rane and Paramesha, 2024; Kenny et al., 2021).
Explanations integrate naturally into risk-managed optimisation:
$\min_{\theta}\ \mathbb{E}_{x,\omega}\!\left[\ell\big(f_{\theta}(x,\omega),y\big)\right]+\lambda_{\mathrm{fair}}\mathcal{L}_{\mathrm{fair}}+\lambda_{\mathrm{xai}}\mathcal{R}\!\left(\mathcal{E}(x,f_{\theta},\omega)\right),$
where $\mathcal{R}$ encodes desiderata such as stability (penalising high $\delta$) and parsimony (penalising $\Omega$). This connects fairness (SPD/DI), calibration (ECE), and logging to the governance criteria C1–C10 used in our evaluation.
Limitation: Even theoretically sound explanations can mislead when they are plausible but unfaithful (e.g., over-reading attention maps), trigger Goodhart effects (Strathern, 1997b) (optimising the appearance of fairness rather than outcomes), or lack distributional validity (narratives drift across domains). Reporting stability, counterfactual checks, and provenance mitigates these risks and makes explanations suitable as audit evidence rather than marketing artefacts. Based on our synthesis findings, there is a dangerous ”Explanation Gap” in how we use these tools. While SHAP (Lundberg and Lee, 2017) and LIME (Ribeiro et al., 2016b) provide useful snapshots, the literature reveals they often create ”plausible but unfaithful” stories. They give us a narrative that makes sense to a human, but because transformers are so complex and non-linear, these stories often don’t match what is actually happening under the hood (Strathern, 1997a). This creates a false sense of security that we think we’ve reached transparency when we’ve really just seen a surface-level summary. Next, we present AI-ready test-beds that computes lifecycle KPIs and logging XAI evidence as audit-ready artifacts.
4.4. AI-ready Test-beds
AI-ready test-beds advance RAI by providing controlled environments to develop, test, and validate AI models with reliability and reproducibility guarantees. In the post-ChatGPT era, such platforms help assess non-determinism, multimodal behavior, and risks like bias, hallucination, and privacy leakage, supporting proportionate assurance and governance alignment. We present representative examples in Table LABEL:tab:ai_testbeds. The key value of these test-beds is in producing audit-ready evidence datasheets/model cards, prompt/tool logs, KPI dashboards, red-team logs, and versioned runs.
Table 7. AI evaluation platforms and initiatives used as de facto test-beds for RAI. We distinguish formal test-beds/platforms from broader labs/initiatives commonly used for evaluation
Name Domain Type RAI-focused features
AI4EU AI-on-Demand Platform (AI4EU, 2019) General AI development Platform Ethical-by-design resources; model/dataset catalogue; support for transparency and explainability workflows (used to prototype responsible AI pipelines).
IEEE Ethical AI Systems Test Bed (IEEE, 2019) Ethical AI development Test-bed/Platform Evaluation against ethical frameworks; human-centred design; fairness-oriented assessment protocols; documentation of decision rationale.
AI Testbed for Trustworthy AI (TNO) (for Applied Scientific Research, 2020) Trustworthy AI Test-bed Robustness and transparency checks; fairness analysis; risk/assurance reporting for deployment contexts.
ETH Zurich Safe AI Lab (Zurich, 2020) Safe and fair AI Lab/Platform Safety-critical validation; robustness/hallucination probing; reliability and interpretability methods for high-stakes settings.
HUMANE AI (HumaneAI, 2020) Human-centric AI Platform/Initiative Alignment with human values; societal-impact and fairness guardrails; guidance for participatory, responsible design.
AI for Good (ITU) ((2017), ITU) Social good / RAI Initiative/Program UN SDG–aligned use cases; ethics and societal-benefit framing; showcases reproducible, responsible deployments.
UKRI Trustworthy Autonomous Systems (TAS) Hub (Research and Innovation, 2020) Autonomous systems Program/Test-bed Accountability and transparency for autonomy; safety cases; conformance/compliance support for real-world systems.
Algorithmic Justice League (Algorithmic Justice League, 2024) Algorithmic fairness Initiative Bias discovery and awareness; audits and mitigation advocacy; widely referenced in fairness evaluation contexts (not a formal test-bed).
AI Commons [^3] https://ai-commons.org/ Open AI collaboration Initiative Open, transparent collaboration patterns; governance-mindful sharing; community evaluation practices (not a formal test-bed).
ClarityNLP (Healthcare) (ClarityNLP, 2021) Healthcare NLP Platform Clinical text pipelines; emphasis on fairness, transparency, and ethical data use; supports domain-specific evaluation.
ToolSandbox (Lu et al., 2024a) Tool-using LLMs Research test-bed Stateful, multi-step/equipped LLM evaluation; probes privacy, fairness, transparency and accountability in interactive tasks.
Chatbot Arena (LMSYS) (Org, 2024) LLM evaluation Community platform Open-source platform where users compare LLMs in head-to-head “arena” battles; pairwise ranking with more than 800k community votes and transparent leaderboards; all evaluation pipelines and data are public for reproducibility.
AI Fairness 360 (AIF360) (Bellamy et al., 2018) Fairness evaluation Toolkit Open-source Python toolkit with 70+ fairness metrics and 9 mitigation algorithms; supports bias detection, fairness explanations and interactive web interface for non-experts.
Zeno (Zeno, 2023) AI evaluation Platform Interactive evaluation platform; enables data exploration across any data type and automatically discovers systematic failures using slice finder; supports interactive charts to compare evaluation metrics across slices and models.
Galileo (Labs, 2023) Generative AI evaluation Platform Automated evaluation metrics for GenAI (RAG and agent metrics plus safety/security metrics); real-time guardrails to block hallucinations, PII and prompt injection; captures corner cases and allows custom metrics.
Arize AI (Dong et al., 2024) AI observability Platform AI observability and LLM evaluation platform that automatically detects model issues, traces root causes via indexed training/validation/production data and fine-tunes performance; includes dashboards for evaluation/performance tracing, explainability and fairness. The Bias Tracing tool monitors fairness metrics (e.g., recall parity, false positive parity, disparate impact) and surfaces features and cohorts contributing to bias.
PromptLayer (PromptLayer, 2022) Prompt engineering Platform Prompt-evaluation platform with visual pipelines for scoring prompts against golden datasets, backtesting with historical data, regression testing and continuous integration of evaluation pipelines; useful for systematic RAG/agent evaluations.
FairX (Sikder et al., 2024) Fairness benchmarking Open-source toolkit Modular benchmarking tool for fairness, data utility and explainability; supports training of fair generative models, evaluation of synthetic data and fairness metrics with pre-, in- and post-processing debiasing techniques.
BEATS bias test suite (Abhishek et al., 2025) LLM bias/ethics evaluation Test suite Benchmark for evaluating bias, ethics, fairness and factuality across 29 metrics for LLM outputs; measures demographic and cognitive biases, ethical reasoning and factuality; aims to diagnose factors driving bias and develop mitigation strategies.
4.4.1. The Lifecycle Evidence Matrix: Mapping Metrics to Engineering Controls
To facilitate the implementation of our proposed evaluation infrastructure, we synthesize the relationship between KPIs, benchmarks, and technical tools. As argued throughout this survey, the effectiveness of Responsible GenAI hinges on the ability to produce portable, verifiable evidence across the lifecycle. Table 8 provides a decision-support mapping that links the KPIs defined in Section 4.2.4 to the specific testbeds and datasets that operationalize them. This matrix enables practitioners to select the appropriate ”tooling stack” based on their specific governance requirements, whether prioritizing fairness audits in Finance or robustness testing in Defense.
Table 8. The ResGenAI Accountability Matrix: Linking KPIs to Benchmarks, Datasets, and Audit Toolkits
KPI Category Primary KPIs Benchmarks & Datasets Audit Toolkits & Testbeds
Fairness & Bias SPD, DI, $\Delta_{\mathrm{acc}}$ SafetyBench (Zhang et al., 2024), HumaniBench (Raza et al., 2025) AIF360 (Bellamy et al., 2018), Fairlearn (Weerts et al., 2023), Arize (Dong et al., 2024)
Safety & Toxicity $E_{h}$ (High-stakes error) HarmBench (Mazeika et al., 2024), SALAD-Bench (Li et al., 2024a) Unitary (Ltd., [n. d.]), Project Moonshot (AI Verify Foundation, 2024), Galileo (Labs, 2023)
Robustness & Security $R$ (Robustness), ASR ALERT (Tedeschi et al., 2024), Rainbow Teaming (Samvelyan et al., 2024) ART (Research, [n. d.]), PromptLayer (PromptLayer, 2022), ToolSandbox (Lu et al., 2024a)
Privacy $P$ (Compliance rate) PrivLM-Bench (Li et al., 2024b), PrivacyLens (Shao and et al., 2024) Privacy Meter (Privacy and Lab, [n. d.]), Google DP (Google, [n. d.]), SecretFlow (Group, [n. d.])
Explainability $E(x)$, LFG (Faithfulness) HELM (Liang et al., 2023), XSTest (Röttger et al., 2024) LIT (Google PAIR, [n. d.]), InterpretML, AIX360 (Arya et al., 2019)
Sustainability $E_{\mathrm{kWh}}$, $C_{\mathrm{CO_{2}e}}$ CodeCarbon (Song et al., 2025) Evidently (Evidently AI, [n. d.]), Arize (Dong et al., 2024), Zeno (Zeno, 2023)
4.4.2. Turning Metrics into Assurance: The ResGenAI Audit Loop
The technical efficacy of ResGenAI remains localized unless it can be translated into portable audit evidence. As visualized in Fig. 10, our framework proposes a three layer loop that bridges the Evaluation Gap identified in Section 6.4. In this cycle, the technical KPIs defined in Section 4.2.4 act as the raw sensory data for the ecosystem. These metrics are packaged into standardized audit artifacts such as model cards and red teaming logs which serve as the portable evidence required by compliance auditors under the EU AI Act (European Union, 2024) or NIST AI RMF (of Standards and, NIST). By establishing this formal “chain of trust”, organizations can move beyond static safety checks toward continuous stewardship where regulatory feedback directly informs the recalibration of technical metrics.
Figure 10. The ResGenAI Audit Loop: translating technical KPIs into portable audit evidence to support regulatory assurance and continuous stewardship across the foundation model lifecycle.
5. Responsible Generative AI Application Across Domains
GenAI systems are increasingly embedded in decision-making processes across diverse domains such as healthcare, finance, human resource, defense and education. While these systems offer significant utility, they also introduce new categories of risks and challenges, as outlined earlier. The challenges and mitigation strategies synthesized in this review are broadly relevant, but their salience and priority vary substantially by sectoral context. We translate the above findings into stakeholder-aligned practices and a deployment-oriented roadmap, with examples and decision pointers. Each subsection below examines domain-specific applications, challenges, and strategies through the lens of our governance-aligned rubric (C1–C10, Section 4.2.1) and lifecycle KPIs (Section 4.2.4). For every domain we (i) identify which rubric criteria are most salient, (ii) assess how the benchmarks in Tables 46 cover those risks, and (iii) indicate which KPIs practitioners should prioritize. Table 9 consolidates this analysis, mapping each sector to its primary criteria gaps and priority KPIs.
5.1. Healthcare
Applications
Generative models improve efficiency in clinical note generation and triage chatbots (Alagha and Ahn, 2023), they raise concerns about accuracy, liability, and bias in medical decision support. Disease detection and drug discovery pipelines (Liu et al., 2023c) illustrate GenAI’s potential for accelerating science, yet demand rigorous validation, explainability, and regulatory alignment before clinical deployment. Synthetic health data helps mitigate privacy risks and dataset imbalance (Walsh et al., 2023), but governance frameworks must ensure that synthetic outputs do not inadvertently reinforce hidden biases. Similarly, tools like Watson for Oncology highlight the need for transparency and interpretability in high-stakes decision-making (Memorial Sloan Kettering Cancer Center, 2024). Even promising breakthroughs, such as GenAI-designed therapeutics, require human oversight on trial integrity, equitable access, and long-term monitoring. In this way, healthcare exemplifies both the promise of GenAI and the critical role of RAI practices, balancing innovation with safety, accountability, and fairness.
ResGenAI challenges
Responsible adoption of GenAI in healthcare is constrained by privacy violations (Shiraz and Lee, 2019), hallucination of critical clinical facts (Li et al., 2023b), algorithmic bias in diagnostics (Obermeyer et al., 2019), unclear accountability for errors (Kaminski, 2022), opacity of generative models (Ghorbani et al., 2019), and difficulty of compliance with traditional regulations (Price and Gerke, 2022). Empirical studies highlight these risks, for example, a 2023 evaluation of 450 consultation transcripts using GPT-4 reported a 1.47% hallucination rate, with 44% of errors classified as major and concentrated in treatment plans (Li et al., 2023b); Rosbach et al. further found a 7% automation-bias error rate in a pathology study where clinicians overturned correct diagnoses by over-trusting AI outputs (Rosbach et al., 2024); and a systematic review of 778 studies showed that explainable AI can both increase and erode clinician trust, underscoring regulatory and accountability challenges (Tjoa and Guan, 2023).
Evaluation context
Mapping these risks to our rubric, the dominant healthcare concerns correspond to C4 (hallucination of clinical facts), C1 (diagnostic bias), and C6 (patient data leakage). Only 7 of 21 benchmarks in Table 4 score above zero on C4 (Misinformation/ Disinformation), and none specifically target medical hallucination. For privacy, only PrivLM-Bench and PrivacyLens provide strong C6 (Privacy) coverage, however, neither addresses HIPAA-specific leakage vectors such as clinical note memorization. C7 (System-level failure) is critical for RAG-based clinical pipelines and is tested by only 4 of 21 benchmarks. The KPIs most relevant to this domain are the high-stakes error rate $E_{h}$ (KPI 9), operationalized as hallucinated clinical facts per $N$ consultations, data privacy compliance $P$ (KPI 2) under HIPAA, and fairness metrics SPD/DI (KPIs 3–4) stratified by patient demographics.
ResGenAI strategies
Several strategies are being employed to enable responsible GenAI adoption in healthcare. Grounding methods (C4, KPI 9) such as RAG can help reduce hallucinations by linking model outputs to verifiable medical sources. Effectiveness should be tracked via $E_{h}$; a reported 1.47% hallucination rate provides a baseline. Human-in-the-loop (HITL) validation (C7, C8, KPI 10) integrates clinician oversight into diagnostic and documentation workflows to mitigate automation bias. Audit frequency $F_{a}$ and resolution time $T_{r}$ should be reported.Privacy-preserving approaches (C6, KPI 2), including differential privacy and federated learning, protect sensitive health records while supporting multi-institutional collaboration. Fairness-aware data augmentation and structured fairness audits (C1, KPIs 3–5) are being applied to improve equity in clinical outcomes. HumaniBench (C1=2 in Table 4) offers a starting template but lacks clinical subgroup stratification. Post-hoc explainability (C9, KPIs 7–8) techniques such as SHAP and LIME are critical for clinician trust.Explanation quality should be assessed via the local faithfulness gap LFG (KPI 8). Alignment with regulatory frameworks (C10) like FDA guidance ensures safe and accountable deployment. Table 6 shows that only HELM and MLCommons v0.5 provide partial transparency support relevant to FDA conformity assessment.
5.2. Finance
Applications
GenAI is transforming multiple facets of finance, with impacts on efficiency, risk management, and personalization (Cao, 2022). Key applications include customer-facing bots that provide interactive banking and wealth-management support, modeling of fraud detection, credit scoring, risk forecasting, and insurance assessment (Cao, 2022), algorithmic trading that uses generative simulations to optimize portfolios and anticipate market shifts (HSBC Research, 2025), personalized financial advice adapted to client risk profiles and long-term goals (Wang et al., 2023b), automated compliance reports and market summaries that accelerate auditing and regulatory filings (Kaminski, 2022), and synthetic financial data generation for stress testing, scenario design, and risk modeling (Cao, 2022). Empirical deployments highlight these advances. For example, PayPal’s fraud detection reduced false positives by 50% in 2023 while processing 10 million transactions daily (Ngai et al., 2011). Charles Schwab’s assistant resolved 62% of client queries and improved satisfaction by 28% in 2023 (Wang et al., 2023b). JPMorgan’s simulator improved trading returns by 12% and reduced value-at-risk by 13% in 2023 (Zhang et al., 2023). HSBC’s AI achieved 95% accuracy in volatility prediction in 2025 (HSBC Research, 2025). Silva et al. also reported RegTech-GPT summaries rated “complete” by 85% of compliance officers compared with 62% baseline (Silva et al., 2024).
ResGenAI challenges
GenAI in finance raises a number of domain-specific risks that complicate its responsible adoption. These include discriminatory lending that replicates historical bias in credit and valuation data , security vulnerabilities where adversarial manipulation targets fraud detection or trading systems, consumer data privacy concerns arising from reliance on sensitive financial records, opacity of GenAI that limits auditability and accountability, hallucination and misleading outputs in customer-facing systems that may distort investment or compliance advice, and market manipulation risk that threatens systemic stability when generative simulations influence trading behavior (Hendry et al., 2022). Empirical use cases underscore these risks. For example, Zillow’s valuation model undervalued homes in minority neighborhoods and reduced appraised equity for underwriting (Adams et al., 2022). Citibank was also fined twenty-five million dollars in 2020 when its flood-zone underwriting model failed auditability standards (Office of the Comptroller of the Currency, 2020).
Evaluation context
The primary finance risks map to C1 (lending discrimination), C3 (adversarial manipulation of trading/fraud systems), C6 (consumer data privacy), and C4 (hallucinated financial advice). No benchmark in Table 4 specifically targets lending discrimination or financial adversarial attacks. The most relevant KPIs are SPD and DI (KPIs 3–4) applied to loan approval rates across demographic groups, robustness $R$ (KPI 6) under financial adversarial scenarios, and audit frequency $F_{a}$ (KPI 10) aligned with SEC and OCC reporting cadences.
ResGenAI strategies
Multiple strategies are being prioritized to address these financial risks. Structured fairness audits (C1, KPIs 3–4) and fairness-aware data augmentation are increasingly required in lending and credit scoring. Post-hoc explainability (C9, KPIs 7–8) methods such as SHAP and LIME support regulators in interrogating opaque credit and trading models. Privacy-preserving techniques (C6, KPI 2), including differential privacy and secure multiparty computation, enable safe collaboration across financial institutions while protecting consumer data. Stronger data provenance frameworks are being developed to track the lineage and validity of synthetic datasets. Alignment with regulatory compliance (C10) frameworks such as Basel III, SEC, and OCC guidelines anchors AI deployments within enforceable legal standards (Kaminski, 2022). Finally, embedding HITL oversight (C7, KPI 10) across lending, trading, and compliance ensures accountability by mandating expert review of generative outputs in high-stakes decisions.
5.3. Education
Applications
GenAI is increasingly shaping critical aspects of education, expanding opportunities for personalization, access, and efficiency (Holmes et al., 2022). Key applications include automated grading that accelerates evaluation and reduces instructor workload, personalized tutoring systems that adapt to student knowledge gaps and learning styles, content generation that produces adaptive curricula and study materials, accessibility tools that deliver multimodal and multilingual support for inclusive education, and interactive simulations that foster experiential and inquiry-driven learning (Chen and Huang, 2023; Lai and Xu, 2024). Recent deployments provide evidence of these advances. Carnegie Learning’s MATHia platform improved algebra test scores by 22% in U.S. high schools in 2023 (Carnegie Learning, 2023). A 2024 Arizona State University pilot found that students using an LLM-powered Socrates chatbot showed 42% higher engagement on surveys (Arizona State University, 2024). A Duolingo study in 2024 reported a 12% increase in learner retention and practice frequency through generative conversational exercises (Duolingo Research, 2024). A U.K. school district pilot in 2024 showed that 87% of students felt classroom chat assistants clarified concepts and 92% of teachers reported time saved (UK Department for Education, 2024).
ResGenAI challenges
Responsible use of GenAI in education faces several pressing concerns. These include cultural insensitivity where models fail to capture local norms and values, student data privacy and particularly minors’ data privacy given the sensitivity of educational records , skills evasion where over-reliance on AI weakens critical thinking and independent problem-solving, and pedagogical opacity that limits educators’ ability to evaluate how generative systems produce outputs (Zawacki-Richter et al., 2019). Empirical use cases underscore these risks. For example, India’s government-run “Newton” homework bot trained only on U.S. and U.K. corpora recommended “pork stew” to rural students in Rajasthan in 2021, violating cultural and nutritional guidelines and prompting suspension of the tool for eleven weeks (UNESCO GEM Report, 2022). A 2023 randomized controlled trial at the University of Cyprus found that physics students allowed to use GPT-4 for weekly assignments scored twenty-three percent lower on conceptual-transfer exam questions despite higher weekly grades, demonstrating over-reliance effects (Pavlou et al., 2023). In 2022, the Utah State Board of Education cancelled a contract for an AI essay-scorer when auditors reported that the vendor could not document the feature vectors or training data, raising concerns about pedagogical opacity and compliance with state education law (Utah State Board of Education, 2022).
Evaluation context
Education risks concentrate on C6 (minors’ data privacy), C4 (hallucinated pedagogical content), C1 (cultural bias), and C7 (system-level failure in tutoring pipelines). No benchmark addresses C6 in the context of minors’ data, and the user inclusivity index $U$ (KPI 11) critical for accessibility in educational settings, lacks any standardized evaluation protocol. Skills-evasion effects (a form of C7 system-level failure) are entirely absent from current safety suites. Priority KPIs include $P$ (KPI 2) under COPPA/FERPA, $E_{h}$ (KPI 9) for pedagogically harmful outputs, and $U$ (KPI 11) for WCAG conformance.
ResGenAI strategies
Multiple strategies are being explored to ensure responsible generative AI in education. Fairness audits (C1, KPIs 3–4) and external ethics audits verify alignment with frameworks such as the UNESCO(UNESCO, 2021a) AI-ethics rubric and can keep bias metrics below accepted thresholds. Privacy-preserving techniques (C6, KPI 2) including homomorphic encryption and federated learning protect minors’ sensitive data while enabling large-scale analysis (Kaissis et al., 2021). Clear regulation and policy guidelines (C10) developed by international and national bodies provide governance and accountability mechanisms. Finally, academic integrity tools (C5, C9) such as watermarking and output tracking are being adopted to prevent plagiarism and monitor AI-assisted work (Zhang and Zhao, 2024).
5.4. Arts and Entertainment
Applications
GenAI is reshaping creative production and cultural participation in the arts and entertainment sector. Prominent applications include creative writing tools that assist in script, story, and lyric development, generative art and design platforms that expand aesthetic experimentation, music composition and production systems that automate melody, arrangement, and mastering, film production support for visual effects, dubbing, and storyboarding, personalized recommendation engines that curate media consumption at scale and virtual influencers and avatars that shape marketing and fan engagement (Crawford, 2021). Recent deployments provide evidence of these advances. For instance, Spotify’s Discover Weekly playlists improved user engagement by 40% in 2023, with subsequent adjustments in 2024 reducing genre bias (Anderson and Lee, 2023). A 2024 marketing campaign featuring AI-driven virtual influencers reached more than ten million users (Crawford, 2021; Traitter and Klein, 2022). Davis et al. released a ten-track AI album in which 62% of listeners in a Spotify trial could not distinguish the music from human compositions (Davis et al., 2024).
ResGenAI challenges
Responsible use of GenAI in arts and entertainment faces a series of ethical, legal, and cultural concerns. These include copyright infringement and intellectual property violation where AI systems reuse creative material without authorization, job displacement that threatens creative labor markets, cultural insensitivity when generative tools fail to respect community values or norms deepfakes and misinformation that erode authenticity and trust, authorship and agency disputes where creative ownership is contested, insufficient consent in data usage for training models, and opaque attribution that obscures the role of AI in the creative process (Schneider and Jörgens, 2022; Elgammal et al., 2019). Empirical use cases underscore these risks. For example, the U.S. Copyright Office’s 2023 ruling in Thaler v. Perlmutter rejected copyright protection for AI-generated works, reinforcing the requirement of human authorship (U.S. Copyright Office, 2023). The public controversy around the 2023 song “Heart on My Sleeve,” which mimicked the voices of Drake and The Weeknd, raised critical questions about derivative works and copyright boundaries. In 2025, WeTransfer updated its terms of service to allow user content for AI training, sparking backlash over transparency and consent (WeTransfer, 2025).
Evaluation context
The dominant risks map to C5 (deepfakes and media integrity), C1 (representational bias in generated content), and C9 (attribution transparency). Only SHIELD and MM-SafetyBench provides explicit C5 evaluation in Table 4, and its scope does not extend to voice cloning or music mimicry, the vectors most salient for this sector. Copyright memorization, a C6-adjacent concern, is not addressed by any privacy benchmark. Priority KPIs are currently underspecified for creative domains; provenance-verification rates and consent-compliance metrics (analogous to KPI 2) are needed but lack standardized protocols.
ResGenAI strategies
Multiple strategies are being adopted to promote responsible use of generative AI in arts and entertainment. Artist collaboration platforms and alternative compensation models are emerging to safeguard livelihoods and redistribute value. Structured bias audits (C1, KPIs 3–4) and cultural sensitivity training are being incorporated into creative AI pipelines to mitigate stereotyping . Deepfake detection tools (C5) are being deployed to maintain authenticity in music, video, and media (Chesney and Citron, 2019). Collective bargaining and union agreements such as those negotiated by SAG-AFTRA establish protections for creative workers (SAG-AFTRA, 2023). International ethical guidelines (C10) and transparency labeling requirements provide standards for disclosure and attribution. Finally, intellectual property management systems (C6, C9) are being advanced to secure consent and compliance. For example, a 2023 copyright detection tool identified AI-generated violations with 95% accuracy (Zirpoli, 2023a; Traitter and Klein, 2022), and these systems require explicit informed consent before creating digital replicas of performances, offering a contractual model for ethical rights management.
5.5. Agriculture
Applications
GenAI is increasingly transforming agricultural practices through precision methods and data-driven innovation. Key applications include precision farming systems that optimize irrigation, fertilization, and pesticide usage, crop yield prediction models that integrate weather, soil, and satellite data, pest and disease detection tools that leverage generative augmentation to improve diagnostic accuracy, virtual agronomy advisors that deliver context-sensitive recommendations to farmers, agricultural robotics for automated harvesting and monitoring, crop breeding simulations that accelerate trait discovery, and weather forecasting for climate adaptation (Costa et al., 2023). While platforms like Microsoft’s FarmBeats (Microsoft Research, 2023) initially achieved significant yield and water-saving gains through IoT and predictive ML , GenAI is now being integrated to augment these systems through generative weather simulation for risk modeling and as LLM-based natural language interfaces . These tools leverage GenAI to translate complex sensor data into actionable, localized advice.
ResGenAI challenges
Responsible adoption of GenAI in agriculture faces several interrelated challenges. These include hallucination and context-inappropriate advice where unreliable outputs may jeopardize crop management decisions, farm data ownership conflicts where smallholders lack control over datasets aggregated by multinational platforms, bias toward major corporations and regions that limits the applicability of tools for smallholder and Global South farmers, overreliance on AI and erosion of local knowledge that risks displacing traditional farming wisdom, IoT security vulnerabilities that expose connected farm systems to cyberattacks, and the persistent digital divide that excludes low-resource communities from advanced AI benefits (Rotz et al., 2019; Sapkota et al., 2024). Empirical use cases underscore these risks. For example, farmers in sub-Saharan Africa reported reduced trust in AI weather advisors when predictions conflicted with lived knowledge, illustrating both hallucination and cultural misalignment. Cases of contested ownership emerged when farm sensor data collected by multinational seed companies were commercialized without equitable farmer compensation (Bronson, 2019). IoT breaches in 2023 European trials highlighted vulnerabilities in sensor networks that exposed proprietary farm-level data.
Evaluation context
Agriculture risks primarily concern C4 (hallucinated agronomic advice), C6 (farm data ownership and privacy), and C7 (system-level failure in IoT-integrated advisory pipelines). No benchmark in our assessment provides geo-aware hallucination testing, and C7 remains untested for sensor-to-advice chains. Priority KPIs are $E_{h}$ (KPI 9) operationalized as harmful agronomic recommendations per $N$ advisory sessions, $P$ (KPI 2) for data governance compliance, and $E_{\text{kWh}}$ (KPI 12) for sustainability reporting—particularly relevant given the energy constraints of low-resource deployment contexts.
RAI strategies
Multiple strategies are being pursued to enable responsible deployment of generative AI in agriculture. Geo-aware retrieval and annotation (C4, KPI 9) techniques are being designed to contextualize model outputs with local weather, soil, and cultural practices, reducing hallucination risks. Data governance contracts (C6, KPI 2) are being negotiated to clarify ownership rights and revenue-sharing models for farm-generated datasets. Embedding human-in-the-loop oversight (C7, KPI 10) into agronomy advisory workflows ensures farmers and extension officers validate AI outputs before implementation. Privacy-preserving IoT integration (C6, KPI 2), including differential privacy and federated learning, protects sensitive farm data across institutions. Finally, alignment with sustainability regulations (C10, KPI 12) such as FAO and EU Common Agricultural Policy frameworks anchors generative AI deployments in enforceable ecological and social standards. Together, these approaches enhance both the equity and safety of generative AI in global agriculture.
5.6. Defense
Applications
GenAI is increasingly influencing defense through simulation, sensing, and decision support. Key applications include novel training and war-gaming environments that generate diverse scenarios and accelerate force development, threat detection and classification across multi-INT streams that fuse imagery signals and text for faster situational awareness, autonomous reconnaissance systems that coordinate teams of uncrewed assets in contested environments, and logistics and supply chain optimization that shortens planning cycles and improves mission readiness (News, 2024). Recent deployments provide evidence of these advances. For instance, DARPA programs such as CODE and successor autonomy efforts demonstrated collaborative uncrewed aircraft operating with minimal communications and supervisor control in flight tests (Defense Advanced Research Projects Agency, 2018). The U.S. Army extended enterprise decision platforms in 2024 to speed analysis and targeting workflows at scale (News, 2024). NATO war-games in 2024 integrated generative scenario simulation and reported that AI-generated vignettes helped surface unanticipated vulnerabilities during exercise analysis.
RAI challenges
Responsible adoption of GenAI in defense introduces ethical operational and legal concerns. These include autonomous weapons and responsibility gaps where assignment of accountability for the use of force remains unresolved, opacity of decision making that limits auditability and traceability in high-stakes contexts, adversarial vulnerabilities that expose perception models and decision aids to spoofing and data poisoning, escalation of conflicts where fast automated responses may narrow decision windows and increase miscalculation risk, ethical and legal gaps regarding compliance with international humanitarian law and rules of engagement, research security concerns over dual-use model capabilities and insider threat, protection of critical infrastructure as AI is integrated into command networks and sensors, and facial recognition deployment where demographic performance differences raise risks in security operations. Empirical use cases underscore these risks. For example, national and alliance exercises reported that AI-supported war-gaming surfaced escalatory flashpoints that required added human oversight (NATO Allied Command Transformation, 2024). Government red-team evaluations and standards work identified numerous novel failure modes in generative systems that could enable misinformation or degrade cyber defences if left unmitigated (National Institute of Standards and Technology, 2025b; WIRED, 2025). Facial recognition evaluations documented demographic differentials that necessitate strict control and testing before operational use (Grother et al., 2019, 2022).
Evaluation context
Defense risks concentrate on C3 (adversarial robustness), C7 (system-level failure in multi-agent autonomy), C8 (red-team realism), and C10 (governance alignment with IHL and NATO principles). C8 is the most critical gap: only 2 of 21 benchmarks (Rainbow Teaming and ALERT) score above zero for adaptive adversary simulation, yet defense applications face the most sophisticated threat actors. C7 is entirely absent for multi-agent autonomous systems. Priority KPIs are robustness $R$ (KPI 6) under military-grade adversarial scenarios, $E_{h}$ (KPI 9) for consequential targeting or escalation errors, and audit frequency $F_{a}$ (KPI 10) aligned with NATO after-action review cadences.
RAI strategies
Multiple strategies are being advanced to ensure responsible deployment. Human in the loop oversight (C7, C8, KPI 10) is prioritized for mission-critical tasks so that trained operators validate model outputs before action. Meaningful human control (C10) is reinforced to keep decisions on the use of force within accountable human authority consistent with international humanitarian law. Explainable AI for defense transparency (C9, KPIs 7–8) adopts traceability and interpretability requirements to support audit and after-action review. Red teaming (C3, C8, KPI 6) and rigorous test and evaluation are expanded to probe adversarial and safety failure modes before fielding. International cooperation (C10) aligns national practices with alliance principles such as NATO’s Principles of Responsible Use and shared assurance processes (NATO, 2024). Ethical review boards and program-level governance embed policy risk assessment data governance and model documentation across the AI life cycle (U.S. Department of Defense, 2024; Defense Innovation Unit, 2020).
5.7. Cross-Applications Synthesis of GenAI Maturity
To provide a critical comparison of how responsible practices are being operationalized across domains, and to connect sectoral risks to the evaluation infrastructure in Sections 4.24.4, we present the extended Synthesis Matrix in Table 9. Beyond risk mechanisms and regulatory drivers, the table now maps each sector to its primary rubric-criteria gaps and priority lifecycle KPIs, making explicit which evaluation evidence is available and where portable audit artifacts are missing.
5.7.1. Stakeholder-Metric Alignment: Operationalizing the Framework
To ensure the sectoral insights in Section 5 translate into actionable governance, we map the relationship between stakeholder roles and our evaluation infrastructure. Based on our synthesized findings, technical frameworks often fail because they do not specify which evidence is required by which actor. In our model, Model Developers utilize the robustness (C3) and toxicity (C2) scores from Table 4 to guide iterative alignment during the fine tuning phase. Conversely, Compliance Auditors and Regulators prioritize the statistical parity (SPD/DI) KPIs (Section 4.2.4) and governance-alignment criteria (C10) to verify adherence to mandates like the EU AI Act (Outeda, 2024) or NIST AI RMF (of Standards and, NIST). Finally, Domain Practitioners in high-stakes fields like Healthcare and Finance rely on the High-Stakes Error Rate ($E_{h}$) and local faithfulness gaps (LFG) to determine the necessary level of human-in-the-loop oversight. By bridging domain-specific mechanisms with normalized technical metrics, our framework provides the ”chain of trust” required to move ResGenAI from theoretical safety to verifiable institutional accountability.
Table 9. Synthesis Matrix: Comparative Analysis of RAI Maturity, Regulatory Pressures, Rubric Criteria Gaps, and Priority KPIs Across Domains.
Sector Dominant Risk Mechanism Technical Control Maturity Primary Regulatory Driver Primary C1 to C10 Gaps Priority KPIs
Healthcare Medical hallucination (Lyu et al., 2025); Diagnostic bias (Gonzalez et al., 2022) Medium: RAG (Wu et al., 2024), HITL, Diff. Privacy (Kaissis et al., 2021) FDA Guidance (U.S. Food and Drug Administration, 2024); HIPAA (Shiraz and Lee, 2019) C4: no clinical hallucination benchmark; C6: no HIPAA specific privacy suite; C7: 4/21 cover system failure $E_{h}$ (9), $P$ (2), SPD/DI (3 to 4), $\Delta_{\text{acc}}$ (5)
Finance Lending discrimination (Mehrabi et al., 2021b); Market manipulation (Hendry et al., 2022) High: Fairness audits (Pfohl et al., 2022); Multi party comp. (Khalid et al., 2023) SEC (U.S. Securities and Exchange Commission, 2021); OCC (Office of the Comptroller of the Currency, 2020) C1: no lending discrimination benchmark; C3: no financial adversarial attack suite SPD/DI (3 to 4), $R$ (6), $F_{a}$ (10)
Education Student privacy (Livingstone et al., 2021); Skills evasion (Pavlou et al., 2023) Low: Watermarking (Jiang et al., 2023); Fed. learning (Yang, 2021) UNESCO AI Rubric (UNESCO, 2021b); State Edu. Laws (Utah State Board of Education, 2022) C6: no minors data protocol; C7: skills evasion untested; KPI 11 lacks standard $P$ (2), $E_{h}$ (9), $U$ (11)
[…]
Agriculture Ownership conflicts (Bronson, 2019); Global South exclusion (Tzachor et al., 2022) Low: Geo aware RAG (Ma et al., 2023); Data contracts (Bronson, 2019) FAO (Costa et al., 2023); EU CAP (Costa et al., 2023) C4: no geo aware hallucination test; C7: IoT advisory pipelines untested $E_{h}$ (9), $P$ (2), $E_{\text{kwh}}$ (12)
Defense Autonomous lethality (International Committee of the Red Cross, 2021); Adversarial spoofing (National Institute of Standards and Technology, 2025a) High: Red teaming (Corporation, 2024); Meaningful HITL (International Committee of the Red Cross, 2021) NATO Principles (NATO, 2024); IHL (United Nations Institute for Disarmament Research, 2025) C8: only 2/21 benchmarks score >>0; C7: absent for multi agent autonomy $R$ (6), $E_{h}$ (9), $F_{a}$ (10)
Legal & Justice Recidivism prediction bias (Skeem and Lowenkamp, 2016); Automated decision opacity (Chesterman, 2021) Low to Med.: Explainability (LIME/SHAP) (Microsoft Research, [n. d.]); Bias audits (Center for Data Science and Public Policy, [n. d.]) EU AI Act Art. 6 (Edwards, 2021); U.S. DOJ Guidelines (Podgor, 2003) C1: no recidivism specific fairness benchmark; C2: legal advice toxicity untested; C10: judicial HITL criteria absent SPD/DI (3 to 4), $E_{h}$ (9), $R$ (6), $F_{a}$ (10)
Manufacturing Predictive maintenance failures (Hashemian, 2010); Adversarial sensor spoofing (Cao et al., 2019) Medium: Digital twins (Mihai et al., 2022); Anomaly detection (Chandola et al., 2009) IEC 62443 (ISO/IEC, 2021b); EU Machinery Reg. (European Union, 2024) C3: no industrial adversarial robustness suite; C4: no sensor drift benchmark; C7: multi robot coordination untested $E_{h}$ (9), $R$ (6), $E_{\text{kWh}}$ (12), $F_{a}$ (10)
6. Discussion
6.1. Synthesis and Interpretation
Our review shows that post-GPT scholarship concentrates on five areas: safety and benchmarking, fairness and governance, explainability and transparency, privacy and security, and sectoral applications. Bias and toxicity receive dense coverage, while deepfakes, sustainability, and system-level failure modes receive relatively less attention. This pattern aligns with the thematic trends we documented and with the growing prominence of multimodal and agentic systems in 2022–2025, which increases the need for evaluation that captures tool use and context.
Closing the accountability chain
To fully answer our primary research question, Who is Responsible?, our synthesis posits that the user must be transitioned from a passive consumer to a systemic guardrail. The current imbalance in GenAI governance, where 74% of organizations use GenAI but only 26% have a formal strategy, often results in the end-user inheriting unmitigated technical risks like PII leakage and hallucination. By triangulating the sectoral failures in Section 5 with the digital literacy gaps identified in Section 3.2.2, we propose a Symmetric Responsibility Model. In this model, developers are accountable for Transparency and Alignment, while users are accountable for Operational Integrity and Dissemination. Digital literacy, therefore, is not merely a skill but a core technical control that must be assessed alongside model-side benchmarks in Section 4.2. Without this user centric governance, even perfectly aligned models remain vulnerable to the last mile failure of human misinterpretation or malicious re contextualization. Table 10 operationalizes this model by mapping specific failure modes to the reciprocal duties of each stakeholder.
Table 10. The Symmetric Responsibility Matrix: Mapping Failure Modes to Stakeholder Duties
Failure Mode Developer Duty (Engineering Controls) User Duty (Operational Integrity)
Hallucination & Misinfo Implement RAG frameworks and calibrated uncertainty reporting (KPI 9) (Wu et al., 2024). Verify factual claims via primary sources and cross modal grounding checks.
PII & Data Leakage Deploy Differential Privacy and automated scrubbing during inference (KPI 2) (Nasr et al., 2023). Practice prompt hygiene by omitting sensitive or proprietary data from inputs.
Adversarial Jailbreaking Conduct continuous red teaming and implement robust system guardrails (C8) (Samvelyan et al., 2024). Adhere to acceptable use policies and report emergent safety bypasses to developers.
Representational Bias Perform fairness audits using intersectional metrics like SPD and DI (KPIs 3 to 4) (Kun, 2015). Critically appraise outputs for stereotyping and contextualize results within local norms.
Over reliance (Automation Bias) Provide model confidence scores and explicit explanation gap warnings (KPI 8) (Strathern, 1997a). Maintain human in the loop (HITL) oversight and preserve independent domain expertise.
Industry Safety Artifacts and Advanced Interpretability.
While academic benchmarks offer standardized metrics, the operational reality of ResGenAI is increasingly defined by safety documentation from frontier labs and emerging mechanistic interpretability techniques, as synthesized in Figure 11. Frontier labs have established documentation protocols as de facto benchmarks for governance facing evidence, directly mapping to our C1 to C10 rubric. OpenAI system cards for GPT 4 and GPT 4o provide granular audits for sycophancy and PII leakage, demonstrating high maturity applications of metric validity through traceable safety refusal evidence. Anthropic utilizes Constitutional AI to automate alignment via RLAIF, outlining tiered safeguards that address malicious actor realism. Meta Llama 3 documentation introduces standardized protocols for multilingual red teaming, vital for addressing cross cultural ethical gaps. Google DeepMind technical reports provide empirical evidence for bias and accuracy trade offs, detailing the overcorrection mechanisms that led to observed representational failures. This industry evidence is complemented by the shift toward mechanistic interpretability, which maps internal neural activations to semantic concepts to offer a more rigorous safety check than behavior level prompts alone. Prominent techniques like Sparse Autoencoders (SAEs) isolate circuits responsible for deceptive or biased reasoning, enabling model steering to suppress harmful features during inference. Integrating these mechanistic audits into the RAI lifecycle addresses the black box challenge by enabling empirical verification of internal safety alignment before deployment.
Figure 11. Integration of industry safety documentation and advanced mechanistic interpretability for governance facing evidence.
ResGenAI sustainability via lifecycle assessment
Sustainability in ResGenAI must transition from qualitative discourse to rigorous quantification grounded in Lifecycle Assessment (LCA) frameworks, as integrated into our governance landscape (Figure 8) and synthesized in Figure 12. This necessitates formalizing carbon reporting and energy benchmarks across the development cycle, including pretraining, fine tuning, and inference. Following our proposed lifecycle KPIs (Section 4.2.4), we advocate for reporting energy consumption and carbon equivalent emissions normalized by throughput, such as per 1K tokens generated. Seminal research indicates that training large scale generative models produces environmental footprints comparable to the lifetime emissions of multiple automobiles, highlighting the need for transparent tracking tools and hardware aware reporting. A sustainable AI ecosystem requires a dual focus on optimizing the computational efficiency of models and leveraging generative insights to advance ecological goals, ensuring that technical innovation supports long term environmental stability.
Figure 12. Quantifying Responsible GenAI Sustainability via Lifecycle Assessment (LCA) in ResGenAI Governance. This framework formalizes carbon reporting and energy benchmarks across the pretraining, fine tuning, and inference stages, enabling a transition to rigorous quantification and supporting a sustainable AI ecosystem.
6.2. Lessons from Industry Deployment Incidents
While our benchmark analysis identifies risk-coverage gaps analytically, real-world deployment incidents from major AI developers offer empirical grounding. We synthesize prominent cases below, mapped to our rubric criteria, to demonstrate how the gaps in Tables 36 manifest in production systems.
Alignment and safety failures (C2, C3, C8) Microsoft Bing Chat integration of GPT-4 in early 2023 produced adversarial persona adoption and emotional manipulation during extended sessions (Roose, 2023), behaviors that emerged through multi-turn prompting, precisely the adaptive adversary pattern captured by C8, for which only 2 of 21 benchmarks scored above zero. OpenAI GPT-4o system card documented sycophantic and manipulative voice-mode outputs 444developers.openai.com/api/docs/models/gpt-4o, illustrating subtler alignment failures beyond toxicity classifiers. Google’s Gemini image generator was suspended in February 2024 after producing historically inaccurate racially diverse depictions of historical figures555blog.google/products/gemini/gemini-image-generation-issue/, demonstrating how overcorrection in bias mitigation (C1) can itself produce fairness failures, an interaction that benchmarks evaluate in isolation.
Hallucination in deployed systems (C4, C9) In Mata v. Avianca (2023), attorneys submitted a brief with six fabricated ChatGPT-generated case citations, resulting in judicial sanctions666caselaw.findlaw.com/court/us-dis-crt-sd-new-yor/2335142.html. These cases underscore that only 7 of 21 benchmarks address misinformation with documented constructs, and that the high-stakes error rate KPI ($E_{h}$, Section 4.2.4) must be operationalized in deployment settings.
Privacy and data leakage (C6) Samsung banned internal ChatGPT use in 2023 after employees inadvertently leaked proprietary source code through conversations (Samsung Electronics, 2023), a usage-path leakage vector that no current privacy benchmark assesses. Nasr et al. (Nasr et al., 2023) demonstrated scalable extraction of memorized training data, including PII, from production ChatGPT via divergence attacks, exceeding the scope of PrivLM-Bench and PrivacyLens.
Deepfake and media integrity (C5) A Hong Kong finance worker transferred approximately \$25 million after a video call with deepfake recreations of company executives (Chen, 2024), and AI-generated robocalls impersonating President Biden targeted the 2024 New Hampshire primary (New Hampshire Department of Justice, 2024). Only SHIELD provides explicit C5 evaluation, and neither incident type, social engineering or democratic interference, falls within its scope.
AI Act enforcement (C10) The Act’s GPAI provisions, effective August 2025, require adversarial testing, incident reporting, and cybersecurity protections for systemic-risk models (European Union, 2024). The European AI Office’s first Code of Practice (European Commission, 2025) operationalizes the transparency and robustness requirements mapped in Table 6. Early compliance experience confirms our finding that organizations struggle to assemble portable evidence across governance obligations, as no single benchmark suite covers more than half of the rubric at the “strong” level.
6.3. Trade-offs between Performance Measures
The scoring formulation offers a transparent and reproducible summary of coverage breadth across ten criteria. It supports comparison within a release and over time when paired with clear documentation of constructs, metrics, and versioned artifacts. We emphasize that the score is a coverage indicator rather than a ranking of overall safety and must be interpreted with the governance crosswalk and with domain risk.
The Fairness-Privacy Paradox: Reconciling Representativeness with the Right to Anonymity
A fundamental tension emerging from our synthesis is the operational conflict between the requirement for Representativeness and the rigid Privacy mandates of the General Data Protection Regulation (GDPR) (European Parliament and Council of the European Union, 2016). While mitigating algorithmic bias (Section 4.2.1) necessitates the collection and processing of high-fidelity, demographically diverse datasets to ensure equitable model performance across protected classes, privacy-preserving standards often mandate data minimization and the obfuscation of sensitive attributes. This creates a structural “Fairness-Privacy Paradox”: without access to granular, protected-class identifiers, developers cannot empirically verify statistical parity or conduct intersectional bias audits (Kun, 2015). Consequently, the legal push for anonymity under GDPR can inadvertently mask the very disparities that ResGenAI frameworks seek to eliminate. We posit that resolving this tension requires a transition from absolute data minimization to the adoption of privacy-enhancing technologies (PETs) such as Differential Privacy and Federated Learning that facilitate rigorous fairness auditing while maintaining individual-level anonymity .
Critical Appraisal of Technical Controls: Synthesis and Reflection
After looking closely at how these systems work in the real world, it’s clear there is a significant gap between what we hope these tools can do and how they actually perform when things get messy. The Explanation-Reliability Gap shows us that while we have many ways to explain model decisions (Dwivedi et al., 2023), these often create stories that sound good but don’t actually match what the model is doing under the hood (Strathern, 1997a). We have found that for big models, these explanations are mostly helpful for developers to find bugs, rather than giving regular people or regulators the true transparency they need.
The Bias-Overcorrection Paradox is another hard lesson we’ve learned from cases like the Gemini incident. When we try too hard to fix fairness using simple keywords, we often end up breaking the model’s ability to be factually accurate or culturally aware. This tells us that fairness isn’t just a number to hit (Kun, 2015), but a deep challenge that needs to change based on where the model is being used. Finally, The Shallow Guardrail Problem reminds us that models can look perfectly safe on paper (Zhang et al., 2024) while still being easy to trick. We’ve seen that models which pass every standard toxicity test can still be pushed into strange or harmful behaviour through basic multi-turn conversations (Roose, 2023). The reality is that our current ”guardrails” often just polish the output without fixing the underlying issues, leaving them vulnerable to the automated attacks we see growing today (Samvelyan et al., 2024).
6.4. Analytical Triangulation: Bridging the Evaluation Gap
The persistent safety failures in production systems are not merely isolated technical oversights; they are symptoms of a systemic “Evaluation Gap” where current benchmarks fail to model the adaptive, iterative pressure applied by real-world adversaries. This section triangulates the theoretical risk taxonomy in Section 3 with the empirical deployment failures documented in Section 6.2 through the lens of our comparative assessment (Section 4.2).
Our assessment in Table 4 reveals a stark deficit in evaluative maturity regarding Malicious Actor Realism (Criterion C8), where only 2 out of 21 surveyed benchmarks: Rainbow Teaming and ALERT provide strong coverage for adaptive pressure. This misalignment directly correlates with the high-profile incidents synthesized in Section 6.2, such as the adversarial persona adoption observed in Bing Chat (Roose, 2023) and the scalable data extraction vulnerabilities reported in ChatGPT (Nasr et al., 2023). These failures underscore that prevailing evaluation suites remain largely static and task-local, leaving them unequipped to measure the non-deterministic tactics encountered in the wild.
Furthermore, the representational failures observed in the Gemini overcorrection incident illustrate a critical breakdown in Governance Alignment (C10), where technical alignment efforts are decoupled from the broader sociotechnical context of deployment. By triangulating these findings, we posit that a ResGenAI operating model must transition from one-shot static probes toward continuous, adaptive evaluation frameworks that mirror the malicious-actor realism and distribution shifts inherent in generative ecosystems.
6.5. Limitations of the Study
Like any survey-based study, this work has several limitations. First, the review follows a pre-specified protocol and focuses primarily on the period from November 2022 to December 2025, with selective inclusion of foundational works for definitional clarity. While this post-ChatGPT window captures the most rapid phase of GenAI deployment, it may exclude emerging studies published after the search cutoff or earlier work that did not explicitly use contemporary terminology. Second, our synthesis relies on publicly available papers, dataset cards, benchmark documentation, and governance artifacts. As a result, proprietary evaluations, internal red-teaming practices, or unpublished safety mitigations may be under-represented. Documentation quality also varies across benchmarks and frameworks, which may affect the completeness of criterion-level scoring.
Third, rubric weights are assigned equally by default, and criterion coverage is determined through structured reading of sources followed by discussion and adjudication in cases of disagreement. Although we conducted sensitivity analyses to assess the robustness of equal-weight assumptions, alternative weighting schemes reflecting sector-specific risk priorities could yield different relative rankings. Fourth, the scoring procedure captures breadth of coverage rather than depth or real-world effectiveness. A benchmark receiving a high rubric score indicates multi-risk awareness and governance alignment, but does not guarantee empirical robustness under adversarial deployment conditions. In this sense, the rubric reflects documented capability rather than independently verified performance.
Fifth, the review synthesizes heterogeneous evidence types, including empirical evaluations, conceptual frameworks, and regulatory analyses. This diversity strengthens thematic coverage but precludes formal meta-analysis or quantitative effect-size aggregation. Our conclusions therefore emphasize structural patterns rather than causal claims. Sixth, governance alignment is assessed at the level of plausible evidentiary support rather than legal compliance. Mapping benchmarks to instruments such as the EU AI Act or NIST AI RMF indicates potential contribution to audit evidence, but does not constitute certification or conformity assessment.
Finally, the rapid evolution of multimodal and agentic systems means that new failure modes may emerge after publication. Continuous updates to evaluation suites, governance frameworks, and deployment practices may shift the relative landscape described here. Replication of our screening and scoring is facilitated by the availability of documented sources, explicit inclusion criteria, and the published rubric; however, interpretive judgment remains inherent in structured qualitative synthesis.
6.6. Research directions
The rubric-based analysis in Section 4.2 indicates that commonly used suites provide strong evidence for toxicity and bias, and more limited coverage for privacy, deepfakes, and system-level failures. As a result, many evaluations remain static and task-local, which limits their usefulness for governance where decision makers need evidence that is portable across audits and versions. Future research should target evaluations that model distribution shifts, adaptive adversaries, and multi-agent coordination with external tools.
For practitioners and policy makers, Tables 3 to 8 map benchmarks to risk families and a governance crosswalk provide a starting point for evidence collection. These tables also reveal gaps that are relevant for conformity assessment under emerging regimes. Applied teams can use the matrix to assemble a minimal set of evaluations that cover privacy, robustness, transparency, logging, and post-market monitoring, and then plan additions for multimodal deepfake risk and adaptive red teaming as required by sectoral context.
We highlight five priorities. First, we need to extend benchmarks to capture system-level failures in tool-using and multi-agent settings, especially around retrieval, planning, and actuation (e.g., API calls, code execution, or robotic steps). Second, fold privacy and provenance checks into routine testing, including membership-inference and training-data extraction probes, plus audits of logging sufficiency. Third, create deepfake risk assessments for multimodal models that cover detection, content provenance, and robustness to evasion. Fourth, shift from static to adaptive evaluation with undisclosed items, targeted red teaming, and tests under distribution shift. Fifth, report sustainability alongside accuracy and robustness energy, emissions, and compute-normalized efficiency to enable compute-proportional evaluation and responsible deployment. These priorities directly address the risk-surface gaps surfaced in this survey.
Teams putting ResGenAI into practice can follow a simple loop. First, scope risks by mapping planned features to common risk families (e.g., bias, toxicity, privacy, and security). Next, select baselines from our matrix so that each release is covered by minimum tests across those areas. Then, add context: turn on logging, enforce dataset/model versioning, and keep lightweight documentation so decisions and results remain traceable. Before launch and later during post-market monitorin, there is need to exercise adaptivity with red teaming and a few undisclosed checks to catch regressions. Finally, track a small set of KPIs across the lifecycle: calibration, error rates in high-stakes flows, audit cadence, exclusivity indicators, and energy/emissions. This routine keeps evaluation aligned with governance expectations and makes evidence easier to reuse when audits occur.
6.7. Systemic Vulnerabilities and Blind Spots in ResGenAI
While GenAI capabilities have accelerated, structural blind spots persist across technical and sociotechnical domains (Kenny et al., 2021). This review identifies several systemic gaps that currently hinder universal assurance:
  • RLHF and Alignment Transparency: Alignment advances, such as RLHF and Constitutional AI, have shifted emphasis from raw capability to controllability. However, the documentation of these alignment pathways is often inconsistent, which complicates the auditing of model refusal behaviour and evidence portability across versions (of Standards and, NIST, NIST).
  • Opaque Proprietary Datasets: Responsible stewardship demands adherence to consent and privacy regimes. However, many frontier models rely on web-scale datasets where provenance, licensing, and curation criteria remain undocumented. Truly representative datasets are often infeasible due to data scarcity or legal limits, leading to persistent data leakage and memorization risks (Nasr et al., 2023).
  • Non-Stationary Safety: Finally, safety is inherently non-stationary; model updates, distribution shifts, and evolving adversarial tactics induce drift. Static benchmarks rarely capture these dynamic failure modes , necessitating a shift toward the continuous monitoring and adaptive evaluation frameworks discussed in Section Section 4.2 .
7. Conclusion
This survey comprehensively synthesized the landscape of ResGenAI by evaluating the alignment between global governance mandates and concrete engineering practices. Our synthesis findings revealed a significant evaluation imbalance in current practices, for example, while technical coverage is mature for surface level AI harms like bias and toxicity, the ecosystem remains acutely vulnerable to privacy leakage, deepfake risks, and the system level failure modes inherent in emergent agentic and tool using environments. Furthermore, we identified that the prevalence of static, task-specific benchmarks severely hinders evidence for institutional audits and regulatory compliance. To bridge this evaluation gap, we proposed a research agenda that prioritizes adaptive, multimodal evaluation and provides a practical stewardship to operationalize technical controls through lifecycle KPIs and standardized audit artifacts. By formalizing the symmetric responsibility model and the ResGenAI audit loop, this work establishes the first integrated framework for bridging technical evaluation with regulatory needs, providing a safe, transparent, and sustainable deployment of GenAI models across domains.
Philosophy and the Way Forward
”Software is eating the world” (Marc Andreessen, 2011), ”AI will eat software” (Jensen Huang, 2017), and perhaps most crucially, ”Philosophy must eat AI. (Schrage and Kiron, 2025)
These shifts capture how technology evolves: from fueling industries, to shaping all digital systems, to raising deep social and ethical questions (McCarthy, 2022). If software rewired markets, AI now rewires trust, knowledge, and decision-making. It cannot remain just a technical pursuit; it must be guided by social values, business needs, and political judgment. We expect The direction of responsible AI will be negotiated not only in research labs but equally in classrooms, communities, and parliaments.
”The measure of intelligence is the ability to change.” — Albert Einstein
This quote encapsulates the core challenge of designing AI systems that can adapt ethically and responsibly to evolving societal norms. Adaptability in AI must go beyond technical performance to include ethical flexibility, ensuring systems align with human values. ResGenAI emphasizes not just technical capabilities but also the ethical imperatives of transparency, fairness, and accountability.
”None of us is as responsible as all of us.” — Authors of this paper
Achieving ResGenAI requires a collective effort among developers, policymakers, and users to mitigate bias, prioritize privacy, and ensure accountability. Collaboration across disciplines and sectors is essential to create equitable and ethical AI systems that serve diverse societal needs.
”For better or worse, benchmarks shape a field.” — David Patterson
Traditional AI benchmarks have prioritized accuracy and efficiency, often neglecting fairness, accountability, and societal impact. Redefining benchmarks to incorporate ethical metrics is critical for guiding the development of AI systems that perform both technically and ethically. The data-driven nature of AI presents persistent challenges, as models often inherit biases and privacy risks from training data. These biases, embedded in AI outputs, require rigorous mitigation strategies to distinguish between useful correlations and harmful patterns. While creating a perfectly unbiased AI system is unattainable due to the evolving nature of societal values, continuous evaluation and adaptation of AI models are essential to maintain alignment with ethical standards.
Human decision-making is influenced by moral reasoning and context, while AI operates strictly within its programmed algorithms. This underscores the importance of fostering AI literacy (Long and Magerko, 2020) among developers and users to ensure responsible deployment and use. To implement ResGenAI effectively, AI systems must prioritize transparency, fairness, and adaptability. They should incorporate ethical reasoning and evolve alongside societal norms. As we advance toward Artificial General Intelligence, the true measure of AI should extend beyond technical proficiency to include ethical adaptability (Blili-Hamelin et al., 2024). The pursuit of ResGenAI is challenging yet indispensable. Ongoing advancements in ResGenAI frameworks, coupled with a growing emphasis on AI ethics, offer a path forward. Prominent voices in the field, including Jensen Huang and Sam Altman, emphasize the need for safe and ethical AI.
8. Acknowledgement
This work received no external funding and was undertaken purely out of a passion for learning and advancing knowledge, and safe and responsible deployment of AI, for the betterment of society.
Appendix A Appendix
A.1. Full Search Strings and Filters
Below we list the precise strings and any database-specific options.
ACM Digital Library Fields: Title/Abstract/Keywords. Years: 2022–2025. Document types: Journal, Proceedings.
("Responsible AI" OR "Ethical AI" OR "AI safety" OR transparency OR accountability
 OR fairness OR "bias mitigation" OR privacy OR equity OR "open science")
AND ("generative AI" OR "large language model*" OR LLM OR LVLM OR "vision-language"
 OR multimodal OR "natural language processing" OR "computer vision")
IEEE Xplore Metadata search (Abstract + Index Terms). Years: 2022–2025. Content type: Journals, Conferences.
(("Responsible AI" OR "Ethical AI" OR "AI safety" OR transparency OR accountability
 OR fairness OR "bias mitigation" OR privacy OR equity) NEAR/5 (LLM OR LVLM OR
 "generative AI" OR multimodal OR "vision-language" OR "large language model*"))
Web of Science Topic search (TS=Title, Abstract, Keywords). Years: 2022–2025.
TS=("Responsible AI" OR "Ethical AI" OR "AI safety" OR fairness OR transparency
 OR accountability OR privacy) AND TS=("generative AI" OR LLM OR LVLM OR multimodal)
arXiv Categories: cs.AI, cs.LG, cs.CV, cs.CL. Date filter applied via API.
(all:"Responsible AI" OR all:"AI safety" OR all:"fairness" OR all:"transparency"
 OR all:"accountability" OR all:"bias mitigation" OR all:"privacy")
AND (all:"generative AI" OR all:"large language model" OR all:"LLM" OR all:"LVLM"
 OR all:"vision-language" OR all:"multimodal")
Scopus Field search (TITLE-ABS-KEY). Years: 2022–2025.
TITLE-ABS-KEY("Responsible AI" OR "Ethical AI" OR "AI safety" OR transparency OR
accountability OR fairness OR "bias mitigation" OR privacy OR equity OR "open science")
AND TITLE-ABS-KEY("generative AI" OR "large language model*" OR LLM OR LVLM OR
"vision-language" OR multimodal OR "natural language processing" OR "computer vision")
A.2. Codebook for rubric C1 to C10
Use Score only explicit tasks and metrics. If a task maps to multiple risks, assign it to the primary stated objective to avoid double counting. Evidence priority Official paper and docs; then dataset or model cards and maintained leaderboards; then widely cited third party reports.
C1 Bias and discrimination
Definition: allocational or representational harms across protected or salient groups in generation or decisions.
Evidence: group labels, group aware metrics, disparity or parity targets, bias stress tests.
Scoring: 0 none; 1 single group axis or indirect proxy; 2 multiple axes with stated metrics and protocols.
C2 Toxicity
Definition: abusive, hateful, profane or harassing content and unsafe advice in outputs.
Evidence: toxicity classifiers or human ratings with thresholds, curated harmful prompt sets.
Scoring: 0 none; 1 one classifier or small probe; 2 multiple probes or raters with thresholds and guidance.
C3 Security and robustness
Definition: failure under adversarial input, perturbations, jailbreaks, or distribution shift.
Evidence: specified threat models, attack recipes, robustness metrics, red teaming logs.
Scoring: 0 none; 1 static or single attack; 2 multiple attacks or shifts with adaptive or transfer evaluation.
C4 Misinformation and disinformation
Definition: factual inaccuracy, persuasion or manipulation, synthetic news or citation fabrication.
Evidence: fact checking tasks, grounded QA, hallucination stress tests, persuasion tests with controls.
Scoring: 0 none; 1 limited fact checks or narrow tasks; 2 diverse tasks with documented constructs and metrics.
C5 Deepfakes and media integrity
Definition: deceptive or synthetic media that undermines provenance or identity.
Evidence: detection tasks, watermark or provenance checks, cross modal authenticity tests.
Scoring: 0 none; 1 single modality or heuristic detector; 2 multi modal detection or provenance evaluation.
C6 Privacy
Definition: leakage of personal or training data through extraction or membership inference.
Evidence: membership inference, extraction benchmarks, memorization audits, privacy accounting.
Scoring: 0 none; 1 one attack or audit without coverage claims; 2 multiple attacks or audits with reporting.
C7 System level failure
Definition: end to end failure in tool use, retrieval augmented generation, or agent pipelines.
Evidence: task graphs, tool calls, rollback or refusal criteria, end to end success and failure logs.
Scoring: 0 model only; 1 limited tool or Retrieval-augmented generation (RAG) probes; 2 end to end tasks with failure and recovery criteria.
C8 Malicious actor realism
Definition: adversary realism through red teaming, iterative prompting, or adaptive pressure.
Evidence: scripted or human in the loop red teams, iterative attack frameworks, coverage reports.
Scoring: 0 none; 1 scripted or static red team; 2 adaptive or iterative red team with metrics.
C9 Metric validity and documentation
Definition: clarity of constructs, metrics, labeling, and reproducibility.
Evidence: task construct definitions, annotation guides, inter rater agreement, confidence intervals.
Scoring: 0 undocumented; 1 partial documentation or single metric; 2 full docs with reliability and uncertainty.
C10 Governance alignment
Definition: traceable mapping to controls or obligations in NIST AI RMF, EU AI Act, ISO or similar.
Evidence: control mappings, records or artifacts usable as evidence, audit trails or benchmark statements.
Scoring: 0 none; 1 informal references; 2 explicit control mapping or evidentiary artifacts.
Appendix B Scoring protocol, reliability, and sensitivity
Procedure Two reviewers independently scored each criterion using this codebook. Evidence was drawn from official papers, public documentation, and widely cited usage reports, following the priority ordering specified in the codebook. Disagreements were resolved through structured discussion and, where needed, adjudication by a third reviewer. The codebook and scoring definitions are given below:
Reliability Scoring followed a two-reviewer protocol with codebook-guided independent assessment (Appendix A.2) and discussion-based adjudication for disagreements. A third reviewer was consulted when consensus could not be reached. This structured qualitative approach prioritizes transparency and traceability of judgments over formal inter-rater statistics, which is appropriate given the interpretive nature of rubric-based benchmark assessment in survey research. The codebook, scoring definitions, and adjudication procedure are documented to enable replication Indices Report the risk coverage index $S_{\mathrm{cov}}$, coverage rate $R$, assurance index $S_{\mathrm{assur}}$, and the overall normalized index $S$ as defined in Section 4.2.1.
Robustness and Sensitivity Analysis To verify that our conclusions do not depend on arbitrary weighting, we performed a Sensitivity Analysis of the composite scoring index S. While the primary results (Table 4) utilize equal weighting across criteria C1 to C10, we recomputed the scores under alternate weights that prioritized high-stakes risks (e.g., C3 Security and C6 Privacy). We observed a high Spearman rank correlation ($\rho>0.85$) between the equal-weight baseline and the weighted models, indicating that the comparative breadth of benchmarks is a stable indicator of evaluative maturity rather than an artifact of the scoring scheme.
Non compliance note The governance crosswalk indicates plausible evidentiary support, not legal compliance.
B.1. Glossary (Key Terms & Acronyms)
Table 1. Glossary of key terms and acronyms used in this survey.
Term / Acronym Category Concise definition
RAI Acronym Responsible AI: fairness, privacy, safety, transparency, accountability across the AI lifecycle.
GenAI Acronym Generative AI: models that synthesize content (text/image/audio/code) from learned distributions.
LLM Acronym Large Language Model (e.g., GPT, Llama): transformer-based models trained on large text corpora.
VLM Acronym Vision–Language Model: multimodal models handling text+image inputs/outputs (e.g., BLIP-2, LLaVA).
LMM Acronym Large Multimodal Model: large-scale models spanning multiple modalities (text, vision, audio).
RAG Acronym Retrieval-Augmented Generation: generation conditioned on retrieved external context.
Agentic AI Term Systems where models plan, select tools, act, and maintain memory; may be multi-agent.
CoT Acronym Chain of Thought: intermediate reasoning steps used to solve tasks.
RLHF Acronym Reinforcement Learning from Human Feedback: preference-based alignment fine-tuning.
PEFT Acronym Parameter-Efficient Fine-Tuning: adaptation with limited trainable parameters (e.g., LoRA).
HITL Acronym Human-in-the-Loop: human oversight/feedback integrated into development or operations.
TEVV Acronym Testing, Evaluation, Verification, and Validation: lifecycle assurance practices.
XAI Acronym Explainable AI: methods to make model behavior intelligible (ante-hoc or post-hoc).
LIME Acronym Local Interpretable Model-agnostic Explanations: local surrogate explanations.
SHAP Acronym SHapley Additive exPlanations: additive attributions based on game theory.
Grad-CAM Term Gradient-weighted Class Activation Mapping: visual saliency for CNN/VLM backbones.
Calibration Term Agreement between predicted confidence and empirical accuracy.
ECE Acronym Expected Calibration Error: bin-wise gap between confidence and accuracy (summary).
Brier score Term Mean squared error of probabilistic predictions vs. outcomes (lower is better).
Bias Term Systematic and unfair deviations across demographic groups, often from unrepresentative data.
Fairness Term Equitable outcomes across protected groups; often assessed via SPD/DI and related metrics.
Equalized odds Term Parity of TPR/FPR across protected groups (conditional on true label).
Robustness Term Performance stability under noise, distribution shift, or adversarial perturbations.
Red teaming Term Structured adversarial testing to elicit failures or unsafe behavior.
Benchmark Term Standardized tasks/metrics for evaluation (e.g., safety, bias, privacy).
KPI Acronym Key Performance Indicator: operational metric tracked across the lifecycle (e.g., SPD, ECE, kWh).
Hallucination Term Outputs that are incorrect, unsupported by evidence, or inconsistent with input/context.
Toxicity Term Harmful, offensive, or unsafe outputs; typically evaluated via specialized benchmarks.
Deepfake Term Synthetic media mimicking real persons/events (face/voice/image/video).
Provenance Term Traceability of data/model/content origin; includes watermarking and content credentials.
System-level failure Term Failures from tool use, planning, memory, or multi-agent interaction.
Governance crosswalk Term Mapping technical evidence (benchmarks/KPIs) to regulatory obligations.
Privacy Term Protecting sensitive data from exposure; operationalized via DP, FL, and other PETs.
DP Acronym Differential Privacy: formal guarantee bounding information leakage about individuals.
FL Acronym Federated Learning: training across clients without centralizing raw data.
PETs Acronym Privacy-Enhancing Technologies: methods to reduce/mitigate privacy risks.
Membership inference Term Attack inferring whether a record was in a model’s training data.
Data extraction Term Recovery of memorized training snippets from a model.
PII Acronym Personally Identifiable Information: data that identifies an individual.
Model card Term Standardized documentation of a model’s intended use, data, metrics, and limits.
Datasheet Term Documentation for datasets (provenance, licensing/consent, collection, intended use).
Post-market monitoring Term Ongoing evaluation and incident response after deployment.
Risk families (C1–C5) Term Rubric: C1 Bias, C2 Toxicity, C3 Security/Robustness, C4 Mis/Disinfo, C5 Deepfakes.
Risk families (C6–C10) Term Rubric: C6 Privacy, C7 System failure, C8 Red-team realism, C9 Metric validity/docs, C10 Governance alignment.
NIST AI RMF Acronym U.S. National Institute of Standards and Technology AI Risk Management Framework.
EU AI Act Acronym EU regulation defining risk tiers and obligations (e.g., transparency, monitoring).
ISO/IEC 42001 Acronym AI management system standard for governance and operational controls.
SAIF Acronym Secure AI Framework (Google): security practices for AI systems.
AI RMF Acronym Artificial Intelligence Risk Management Framework (shorthand for NIST AI RMF).
GDPR Acronym General Data Protection Regulation (EU).
HIPAA Acronym Health Insurance Portability and Accountability Act (USA).
FDA Acronym Food and Drug Administration (USA).
OCC Acronym Office of the Comptroller of the Currency (USA).
SEC Acronym Securities and Exchange Commission (USA).
PRISMA Acronym Preferred Reporting Items for Systematic Reviews and Meta-Analyses.
SOTA Acronym State-of-the-Art.
IHL Acronym International Humanitarian Law.
NIST Acronym National Institute of Standards and Technology.
NATO Acronym North Atlantic Treaty Organization.
UNESCO Acronym UN Educational, Scientific and Cultural Organization.
SAG-AFTRA Acronym Screen Actors Guild–American Federation of Television and Radio Artists.
WGA Acronym Writers Guild of America.
AIF360 Acronym IBM AI Fairness 360: fairness metrics and mitigation toolkit.
Fairlearn Acronym Microsoft toolkit for fairness assessment and mitigation.
Table 2. Core synthesis of the literature review: theme, citation, year, and focus.
Theme Citation Year Focus
Safety/Benchmarks (Liu et al., 2024) 2024 Multimodal benchmark to assess safety risks in vision–language models.
Safety/Benchmarks (Zhang et al., 2024) 2024 General safety benchmark for LLMs across risk categories.
Safety/Benchmarks (Gehman et al., 2020) 2020 Toxic degeneration measurement using curated prompts to stress test LMs.
Safety/Benchmarks (Li et al., 2024a) 2024 Hierarchical, comprehensive safety benchmark for language models.
Safety/Benchmarks (Mazeika et al., 2024) 2024 Standardized evaluation for red teaming and refusal robustness.
Safety/Benchmarks (Samvelyan et al., 2024) 2024 Diverse adversarial prompt generation for automated safety stress testing.
Safety/Benchmarks (Wang et al., 2023a) 2023 Broad audit of trustworthiness, including safety and robustness, for GPT models.
Safety/Benchmarks (Tedeschi et al., 2024) 2024 Benchmarks red teaming effectiveness for assessing safety responses.
Fairness/Governance (Mittal et al., 2024) 2024 Dataset design practices aligned with fairness, privacy, and regulatory norms.
Fairness/Governance (Rodolfa et al., 2021) 2021 Minimal fairness–accuracy tradeoff in public policy ML tasks.
Fairness/Governance (Chouldechova, 2017) 2017 Disparate impact and fairness criteria in recidivism risk prediction.
Fairness/Governance (Barocas and Selbst, 2016) 2016 How big-data practices can create disparate impact and legal risk.
Fairness/Governance (Pessach and Shmueli, 2022) 2022 Survey of algorithmic fairness methods and evaluation across domains.
Fairness/Governance (Kaur et al., 2022b) 2022 Trustworthy AI principles and technical enablers for deployment.
Fairness/Governance (Dignum, 2019) 2019 Responsible AI framework grounded in ethics and governance.
Fairness/Governance (Laux et al., 2024) 2024 Trustworthiness concepts in relation to EU AI Act compliance.
Explainability/XAI (Ribeiro et al., 2016b) 2016 LIME: model-agnostic local explanations of predictions.
Explainability/XAI (Selvaraju et al., 2017) 2017 Grad-CAM for visual explanations in CNNs.
Explainability/XAI (Arrieta et al., 2020) 2020 Concepts, taxonomies, and challenges in explainable AI.
Explainability/XAI (Dwivedi et al., 2023) 2023 Review of XAI ideas/techniques with practical guidance.
Explainability/XAI (Saeed and Omlin, 2023) 2023 Meta-survey of XAI challenges and future directions.
Explainability/XAI (van der Waa et al., 2021) 2021 Rule-based vs. example-based explanations: fidelity and user value.
Explainability/XAI (Mosca et al., 2022) 2022 Review of SHAP-based explanation methods for NLP.
Explainability/XAI (Nauta et al., 2023) 2023 Systematizes quantitative evaluation protocols for explanations.
Privacy/Security (Dwork, 2006) 2006 Differential privacy as a formal framework for privacy protection.
Privacy/Security (McMahan et al., 2017) 2017 Federated learning for decentralized training with privacy benefits.
Privacy/Security (Nasr et al., 2023) 2023 Large-scale extraction of memorized training data from LMs.
Privacy/Security (Mattern et al., 2023) 2023 Membership inference via neighborhood comparison against LMs.
Privacy/Security (He et al., 2025a) 2025 Label-only membership inference attacks for pre-trained LMs.
Privacy/Security (Li et al., 2024c) 2024 Benchmark for privacy risks in LMs at multiple levels.
Privacy/Security (Mireshghallah et al., 2024) 2024 Privacy leakage tests using contextual-integrity evaluations.
Privacy/Security (Staab et al., 2024) 2024 Assesses large models for anonymization and de-identification tasks.
LLM Safety (Bai et al., 2022) 2022 Constitutional AI alignment to reduce harmful outputs.
LLM Safety (Weidinger et al., 2021) 2021 Maps ethical/social risks of LMs and mitigation avenues.
LLM Safety (Huang et al., 2024b) 2024 Positions trustworthiness dimensions and evaluation needs for LMs.
LLM Safety (Lin et al., 2024) 2024 Survey of debiasing and hallucination-reduction techniques.
LLM Safety (Wang et al., 2023c) 2023 ChatGPT applications, limitations, risks, and safeguards overview.